### v2

# Testing Most Influential Sets

Abstract
Small subsets of data with disproportionate influence on model outcomes can have dramatic impacts on conclusions, with a few data points sometimes overturning key findings. While recent work has developed methods to identify these most influential sets, no formal theory exists to determine when their influence reflects genuine problems rather than natural sampling variation. We address this gap by developing a principled framework for assessing the statistical significance of most influential sets. Our theoretical results characterize the extreme value distributions of maximal influence and enable rigorous hypothesis tests for excessive influence, replacing current ad-hoc sensitivity checks. We demonstrate the practical value of our approach through applications across economics, biology, and machine learning benchmarks.
1 Introduction
Machine learning (ML) models and statistical inferences can be highly sensitive to small subsets of data. In many applications, just a handful of samples can overturn key conclusions: two countries nullify the estimated effect of geography on development (Kuschnig et al., 2021), a single outlier flips the sign of a treatment effect (Broderick et al., 2021), or a small group of individuals drives disparate outcomes in algorithmic decision-making (Black & Fredrikson, 2021). These most influential sets — data subsets with the greatest influence on model predictions — are central to questions of interpretability, fairness, and robustness in modern machine learning (see, e.g., Black & Fredrikson, 2021; Chen et al., 2018; Chhabra et al., 2023; Ghorbani & Zou, 2019; Sattigeri et al., 2022).
Despite their practical importance, practitioners lack principled tools to assess whether a set’s influence is genuinely problematic. Current practice relies on domain expertise and ad-hoc sensitivity checks, while approximate methods such as influence functions (Koh & Liang, 2017; Fisher et al., 2023; Schioppa et al., 2023) systematically underestimate the impacts of sets and extreme cases (Basu et al., 2020; Koh et al., 2019). Recent work highlights both the promise and challenges of most influential subsets — small sets can drive results even in randomized trials (Broderick et al., 2021; Kuschnig et al., 2021), heuristic algorithms can fail in simple settings (Hu et al., 2024; Huang et al., 2025), and influence bounds remain an active area of research (Moitra & Rohatgi, 2023; Freund & Hopkins, 2023; Rubinstein & Hopkins, 2024). What remains missing is a principled method to distinguish natural sampling variation from genuinely excessive influence.
We develop a statistical framework for assessing the significance of most influential sets. By focusing on linear regression — a tractable, interpretable, and widely-used setting that underlies many modern methods (Rudin, 2019) — we derive the exact asymptotic distributions of maximal influence. We show that two distinct regimes emerge depending on the size of the influential set: when the size is fixed, maximal influence converges to a heavy-tailed Fréchet distribution; when the size grows with the sample, maximal influence converges to a well-behaved Gumbel distribution. Our results enable principled hypothesis tests for excessive influence, replacing ad-hoc diagnostics with rigorous statistical procedures. We demonstrate their practical value via applications across economics, biology, and machine learning benchmarks, resolving ambiguous cases where influential sets drive contested findings.
1.1 Contributions
We present a comprehensive analysis of the influence of most influential sets, both theoretically and in practical applications. To summarize, our main contributions are:
  1. 1.
    Theoretical foundations. We derive distributions for the influence of most influential sets, establishing their extreme value behavior and enabling statistical testing.
  2. 2.
    Efficient implementation. We provide computationally efficient procedures for evaluating influence, making our approach practical for real-world applications.
  3. 3.
    Empirical validation. We demonstrate the utility of our framework across domains, resolving the contested “Blessing of Bad Geography” in economics, assessing robustness in biological data of sparrow morphology, and auditing fairness in ML benchmark datasets.
1.2 Outline
The remainder of the paper is structured as follows. Section 2 introduces the problem of most influential sets and formalizes the setting. Section 3 presents our theoretical results on the distribution of maximal influence. Section 4 demonstrates the practical merits of our framework through simulations and empirical applications. Section 5 discusses implications, limitations, and future directions, and Section 6 concludes.
2 Preliminaries and Background
Practitioners routinely encounter situations where small subsets of data points drive key conclusions. Consider the following scenarios:
  • Scientific discovery: Rugged terrain generally hinders economic development, but not in Africa. What if this striking result is driven by just two small island nations?
  • Fairness auditing: An algorithmic decision-making system produces vastly different outcomes for a protected group. What if this disparity is explained by only a handful of data points?
  • Data cleaning: A single influential data point among a thousand samples flips a strong correlation to a null result. Should we trust the original finding or the one without the outlier?
  • Data preprocessing: A microcredit experiment shows negligible outcome variations overall, except for a few outliers. How should we prepare and analyze the sample?
At the core of these examples lie most influential sets, which exert disproportionate influence on an estimate or prediction. These sets are intuitive to interpret, directly tied to the quantity of interest, and provide a new dimension for assessing estimates by highlighting their support in the data.
2.1 Formal Problem Statement
We consider a supervised learning task with input space $\mathcal{X}\subset\mathbb{R}^{P}$ and target space $\mathcal{Y}\subset\mathbb{R}.$ The goal is to learn a function $f(\theta,\cdot):\mathcal{X}\mapsto\mathcal{Y}$ parameterized by $\theta\in\mathbb{R}^{Q}.$ Given training data $\left\{\left(x_{n},y_{n}\right)\right\}_{n=1}^{N}$ and a loss function $\mathcal{L}\left(\cdot,\cdot\right),$ we learn parameters by solving
$\hat{\theta}=\underset{\theta\in\mathbb{R}^{Q}}{\arg\min}\sum_{n=1}^{N}\mathcal{L}\left(f(\theta,x_{n}),y_{n}\right).$
Let $\left[N\right]=\left\{1,\dots,N\right\}$ and denote both an index set and its corresponding subsample as ${\mathbb{S}}\subset\left[N\right].$ For any subset $\mathbb{S}$}, we use a subscript $\hat{\theta}_{-{\mathbb{S}}}$ to denote a quantity $\theta$ without $\mathbb{S}$}, i.e.
$\hat{\theta}_{-{\mathbb{S}}}=\underset{\theta\in\mathbb{R}^{Q}}{\arg\min}\sum_{n\not\in{\mathbb{S}}}\mathcal{L}\left(f(\theta,x_{n}),y_{n}\right).$
Definition (Most Influential Set).
For a positive integer $k\ll N,$ the $k$-most influential subset is
${\mathbb{S}}_{k}^{\max}\coloneqq\underset{{\mathbb{S}}\subset\left[N\right],\lvert{\mathbb{S}}\rvert\leqslant k}{\arg\max}\Delta\left({\mathbb{S}};\phi\right),$
where $\Delta\left({\mathbb{S}};\phi\right)=\phi(\hat{\theta})-\phi(\hat{\theta}_{-{\mathbb{S}}})$ is the influence of subset $\mathbb{S}$} on target function $\phi:\mathbb{R}^{Q}\mapsto\mathbb{R}.$ The maximum influence is denoted as $\Delta^{\max}=\Delta\left({\mathbb{S}}_{k}^{\max};\phi\right).$
Research Question.
What is the probability distribution of $\Delta^{\max},$ and how can we use it to distinguish excessive influence from natural sampling variation?
2.2 Influence Functions vs. Exact Influence
A common approach to study influence is via influence functions (Fisher et al., 2023). These are motivated by reweighing via the perturbation
$\hat{\theta}(\epsilon;{\mathbb{S}})\coloneqq\underset{\theta\in\mathbb{R}^{Q}}{\arg\min}\frac{1}{N}\sum_{n\not\in{\mathbb{S}}}\mathcal{L}\left(f(\theta,x_{n}),y_{n}\right)+\epsilon\sum_{i\in{\mathbb{S}}}\mathcal{L}\left(f(\theta,x_{i}),y_{i}\right).$
Setting $\epsilon=0$ recovers $\hat{\theta}$, while $\epsilon=-N^{-1}$ yields $\hat{\theta}_{-{\mathbb{S}}}.$ The influence function is the linear approximation at $\epsilon=0$:
$\hat{\theta}(\epsilon;{\mathbb{S}})\approx\mathcal{I}({\mathbb{S}})\coloneqq\frac{\mathrm{d}\hat{\theta}(\epsilon;{\mathbb{S}})}{\mathrm{d}\epsilon}\bigg|_{\epsilon=0}.$ (1)
This yields the following first-order estimate of influence:
$\Delta\left({\mathbb{S}};\phi\right)\approx-N^{-1}\frac{\mathrm{d}\phi\left(\hat{\theta}(\epsilon;{\mathbb{S}})\right)}{\mathrm{d}\epsilon}\bigg|_{\epsilon=0}.$
While influence functions are computationally convenient, they are unreliable even for simple models (Basu et al., 2020; Hu et al., 2024; Huang et al., 2025; Koh et al., 2019). In particular, they systematically underestimate the impact of (a) sets of data points and (b) highly influential data points. This occurs because the first-order approximation cannot reflect higher-order effects from the interplay between data points or differential leverage scores.
Exact Influence
In this work, we focus on the exact influence of subsets in a tractable but ubiquitous setting. We avoid linear approximations to accurately assess the most influential sets of interest, where extreme behavior dominates and first-order approximations fail most dramatically.
3 Proposed Approach
Consider the standard linear regression model, where $f$ is a linear function relating features $X\in\mathbb{R}^{P}$ to the outcome $Y\in\mathbb{R}.$ Stacking the observed training sample yields the design matrix $\mathbf{X}\in\mathbb{R}^{N\times P}$ and outcome vector $\mathbf{y}\in\mathbb{R}^{N}.$ We assume that $\mathbf{X}^{\prime}\mathbf{X}$ is invertible and remains so after removing any subset. [^1] Alternatively, one may consider the penalized estimator with $\mathbf{X}^{\prime}\mathbf{X}+\lambda\mathbf{I},$ where $\lambda>0$ creates a ridge that guarantees invertibility. Let $\mathbf{H}=\mathbf{X}\left(\mathbf{X}^{\prime}\mathbf{X}\right)^{-1}\mathbf{X}^{\prime}$ denote the hat matrix, with leverage scores along its diagonal $\mathbf{h}$. The ordinary least squares (OLS) estimator is
$\hat{\theta}=\underset{\theta}{\arg\min}\lVert Y-X\theta\rVert^{2}=\left(\mathbf{X}^{\prime}\mathbf{X}\right)^{-1}\mathbf{X}^{\prime}\mathbf{y},$
with predictions $\hat{\mathbf{y}}=\mathbf{X}\hat{\theta}$ and residuals $\mathbf{r}=\mathbf{y}-\hat{\mathbf{y}}$.
We focus on the influence of subsets on a particular coefficient of interest for interpretation. For simplicity, we assume a univariate model with a positive coefficient, orthogonalize features where necessary, and set the target function to $\phi(\theta)=\theta_{1}.$
3.1 Exact Influence Formulas
The influence of a single observation $i$ is well-known (Belsley et al., 1980; Cook, 1979) to be
$\Delta\left(\{i\}\right)=\frac{x_{i}r_{i}}{\sum_{n}x_{n}^{2}}\frac{1}{1-h_{i}},$ (2)
where $h_{i}$ is the leverage score and $r_{i}$ the residual of observation $i.$
For sets of observations, we can derive a closed-form expression via recursive application of Equation 2 and updating formulae for $r_{i}$ and $h_{i}.$ (Details are provided in the Appendix.) The following closed form solution is particularly convenient:
Proposition 1.
The influence of set $\mathbb{S}$} on $\hat{\theta}$ is
$\Delta\left({\mathbb{S}}\right)=\frac{\sum_{i\in{\mathbb{S}}}x_{i}r_{i}}{\sum_{n\not\in{\mathbb{S}}}x_{n}^{2}}.$ (3)
Proof sketch.
For a single observation, one can show that $\Delta\left(i\right)=\frac{x_{i}r_{i}}{\sum_{n\neq i}x_{n}^{2}}.$ Then, let ${\mathbb{S}}=\left\{1,2\right\}$ and define $D\coloneqq\sum_{n}x_{n}^{2}$ for simplicity. Evidently,
$\Delta\left(\{1,2\}\right)$ $=\frac{x_{1}r_{1}}{D_{-\{1,2\}}+x_{2}^{2}}+\frac{x_{2}\left(r_{2}+x_{2}\Delta\left(\{1\}\right)\right)}{D_{-\{1,2\}}}$
$=\frac{x_{1}r_{1}}{D_{-\{1,2\}}+x_{2}^{2}}+\frac{x_{2}r_{2}}{D_{-\{1,2\}}}+\frac{x_{2}^{2}\left(x_{1}r_{1}\right)}{D_{-\{1,2\}}\left(D_{-\{1,2\}}+x_{2}^{2}\right)}$
$=\frac{\left(D_{-\{1,2\}}+x_{2}^{2}\right)x_{1}r_{1}}{D_{-\{1,2\}}\left(D_{-\{1,2\}}+x_{2}^{2}\right)}+\frac{x_{2}r_{2}}{D_{-\{1,2\}}}=\frac{x_{1}r_{1}+x_{2}r_{2}}{D_{-\{1,2\}}},$
where the second term in the first line corrects $r_{2}$ to reflect the removal of observation $1,$ which the second line expands. The third line merges terms one and three by transforming to a common denominator, and the fourth line simplifies the expression. Assuming this identity holds for $\lvert{\mathbb{S}}\rvert=K$ we can show by induction that it holds for $\lvert{\mathbb{S}}\rvert=K+1,$ and the result follows. Details are provided in the Appendix. ∎
Proposition 1 elegantly reveals the additive structure of individual contributions in the numerator and the multiplicative adjustment from the remaining data in the denominator. This representation enables efficient computation without explicitly forming leverage scores for each subset, making our approach computationally tractable for large datasets.
3.2 Extreme value distribution
We now turn to the distribution of $\Delta({\mathbb{S}})$ for the most influential set, ${\mathbb{S}}_{k}^{\max}.$ Since this quantity is defined by an extremal operation (maximization over all possible subsets), its asymptotic behavior is governed by extreme value theory. Specifically, we seek the limiting extreme value distribution (EVD) $H$ such that $\Delta^{\max}\in\mathrm{MDA}\left(H\right),$ i.e., $\Delta^{\max}$ lies in the maximum domain of attraction of $H.$
Two canonical EVDs are of particular interest: the Fréchet (Type II) distribution $\Phi_{\alpha}$ for heavy-tailed variables, and the Gumbel (Type I) distribution $\Lambda$ for light-tailed variables. We distinguish two practically relevant regimes based on how the subset size $k$ scales with sample size $N$:
  1. 1.
    Constant-size sets: $k$ remains fixed as $N\to\infty.$
  2. 2.
    Relative-size sets: $k$ grows proportionally with $N,$ i.e., $k=pN$ for some $p\in(0,1).$
Both regimes appear in practical applications (see, e.g, Broderick et al., 2021; Kuschnig et al., 2021), but they yield fundamentally different asymptotic behavior with important implications for significance testing.
3.2.1 Constant-size Sets
Theorem 1 (EVD for constant-size sets).
Suppose $\mathbb{E}\left[X^{2}\right]<\infty,$ and that $X_{i},R_{i}$ have polynomial tails with coefficients $\xi_{x},\xi_{r}<\infty$. If $\lvert{\mathbb{S}}_{k}^{\max}\rvert$ remains constant as $N\to\infty,$ then
$\lim_{N\to\infty}\Delta^{\max}\sim\text{Fréchet}(a,b,\xi),$
with location parameter $a,$ scale parameter $b,$ and shape parameter $\xi=\min\{\xi_{x},\xi_{r}\}.$
Proof sketch.
Let $C\coloneqq\sum_{i\in{\mathbb{S}}}X_{i}R_{i}$ and $D\coloneqq\sum_{n=1}^{N}X_{n}^{2}.$ Notice that $C$ and $D^{-1}_{-{\mathbb{S}}}$ are asymptotically independent. Since $X_{i}$ and $R_{i}$ have polynomial tails with coefficients $\xi_{x},\xi_{r},$ their product satisfies $C\in\mathrm{MDA}(\Phi_{\xi}),$ with $\xi=\min\{\xi_{x},\xi_{r}\},$ and its upper tail behaves like the tail of $\max\{X_{i}R_{i}\}$ for $i\in{\mathbb{S}}_{k}^{\max}.$ Lemma 1 shows that the inverse sum $D^{-1}_{-{\mathbb{S}}}\in\mathrm{MDA}(\Lambda)$, and the product $CD^{-1}_{-{\mathbb{S}}}$ inherits the Fréchet behavior from $C$ by Lemma 2. See the Appendix for details. ∎
This result shows that for constant-size sets, $\Delta^{\max}$ exhibits heavy-tailed Fréchet behavior, implying that even a few observations can exert extreme influence with non-negligible probability.
Corollary 1.
If the tail coefficients of both $X_{i}$ and $R_{i}$ is infinite, then
$\lim_{N\to\infty}\Delta^{\max}\sim\text{Gumbel}(a,b).$
3.2.2 Relative-size Sets
When the most influential set grows proportionally with the sample size, the central limit theorem dominates the asymptotic behavior:
Theorem 2 (EVD for relative-size sets).
If $\{X_{n}R_{n}\}_{n=1}^{N}$ satisfies the conditions of a central limit theorem (CLT) and $\lvert{\mathbb{S}}_{k}^{\max}\rvert$ grows proportionally with $N,$ then
$\lim_{N\to\infty}\Delta^{\max}\sim\text{Gumbel}(a,b).$
Proof sketch.
When $\lvert{\mathbb{S}}_{k}^{\max}\rvert=pN,$ for $p\ll 1$, the numerator $C$ grows at the rate $\mathcal{O}(N).$ By the CLT, $C/\sqrt{N}\sim\mathcal{N}(\mu,\sigma^{2})$ as $N\to\infty$. Hence, the product $CD^{-1}_{-{\mathbb{S}}}$ lies in the maximum domain of attraction of the Gumbel distribution, following Lemma 2 and Corollary 2. See the Appendix for details. ∎
For relative-size sets, $\Delta^{\max}$ converges to a well-behaved Gumbel distribution with exponentially decaying tails, in contrast to the heavy-tailed Fréchet behavior of constant-size sets. This result holds regardless of the underlying distributions of $X$ and $R$ as long as the variance of $X_{i}\cdot R_{i}$ is finite.
3.3 Implementation and Computation
With theoretical results established, we turn towards practical implementation. Our procedure follows three steps:
  1. 1.
    Determine the relevant extreme value distribution. We select between the Gumbel and Fréchet families based on the hypothesized set size and the tail behavior of $X$ and $R$. We estimate tail coefficients using maximum likelihood estimation (MLE; Smith, 1985; Bücher & Segers, 2017). If $1/\max\{\xi_{x},\xi_{r}\}$ is sufficiently close to zero, we default to the Gumbel distribution (per Corollary 1 and Theorem 2). Otherwise, we use the Fréchet distribution with shape parameter $\xi=\max\{\xi_{x},\xi_{r}\},$ following Theorem 1.222Small values of $\xi$ correspond to extremely heavy-tailed distributions where the variance ($\xi\leqslant 2$) or even the mean ($\xi\leqslant 1$) become infinite. Such extreme cases pose practical challenges for statistical inference.
  2. 2.
    Estimate location and scale parameters. We estimate the location and scale parameters $a,b$ using the block maxima method (Coles, 2001; De Haan & Ferreira, 2006). We divide the sample (excluding ${\mathbb{S}}_{k}^{\max}$ for robustness) into $M$ blocks of size $N/M$, compute $\Delta^{\max}$ for each block, and use MLE based on these draws.
    Since selecting the maximum out of $N/M$ observations reduces the expected maximum compared to the full sample, in the Gumbel case a bias correction can be applied. More specifically, we know that
    $F^{N}(x)\xrightarrow{d}\text{Gumbel}(a,b)\quad\text{and}\quad[F^{N/M}(x)]^{M}\xrightarrow{d}\text{Gumbel}(a,b),$
    which yields the location correction $\tilde{a}=\hat{a}+b\log(M)$, where $\hat{a}$ is the MLE.
  3. 3.
    Perform hypothesis testing. We test the null hypothesis $H_{0}$ that the observed influence reflects natural sampling variation against the $H_{1}$ of excessive influence. Using the estimated parameters, we compute the $p$-value as $P(\Delta^{\max}\geqslant\delta_{\text{obs}})$ where $\delta_{\text{obs}}$ is the observed maximum influence.
Computational Efficiency.
Our procedure is computationally convenient, allowing for application to large and varied datasets. The maximum likelihood steps are simple and well-behaved, optimizing over only two parameters in the Gumbel case. The primary computational constraint stems from finding most influential sets — we need to approximate $\Delta^{\max}$ for the $M$ block maxima estimates (Price et al., 2022). For computational tractability, we use an adaptive greedy algorithm (Hu et al., 2024; Kuschnig et al., 2021) with complexity $\mathcal{O}(Mk)$ and considerably reduced runtime from our closed-form influence formula for sets in Proposition 1.
4 Experiments
We validate our theoretical predictions and demonstrate practical utility through controlled simulations and real-world applications spanning economics, biology, and machine learning.
4.1 Simulation Study
Figure 1 illustrates our approach on a simple linear regression with one moderately influential point due to high leverage. Panel A visualizes the data, significance thresholds (at the 10, 5, and 1% significance levels) as a function of predictor and response values. Panel B presents the underlying extreme value analysis: block maxima inform the estimated Gumbel distribution, yielding a $p$-value of 0.04 for the observation of interest.
Figure 1: Illustration of our methodology on a simple linear regression with a moderately influential observation. Panel A depicts observations, estimated regression lines with and without the influential point, and conditional significance regions at the 10, 5, and 1% levels (dotted lines). Panel B shows the extreme value analysis: a histogram of block maxima, fitted Gumbel distribution with (solid) and without (dashed) bias correction, and the resulting $p$-value for the observation of interest.
4.1.1 Convergence to Extreme Value Distributions
We verify that maximal influence converges to the predicted extreme value distributions. Across for scenarios combining standard Normal and $t(5)$ distributions for $X,R,$ we simulate $1,000$ datasets of sizes $N\in\{100,200,500,1000,2000\}.$
Table 1 shows inverse shape estimates $\hat{\xi}^{-1}$ for the setup with $N=100.$ As predicted by theory, the Normal-Normal case yields clear Gumbel behavior ($\xi^{-1}=0$) for all sample sizes (see Table A1). The heavy-tailed cases exhibit predicted Fréchet behavior, with shape estimates matching theoretical predictions ($\xi^{-1}=0.2$), confirming the the finite sample applicability of Theorem 1. Particularly the convergence of the t(5)t(5)–Normal case has slower convergence rates to the limiting distribution due to the instability of $D^{-1}_{\mathbb{S}}$ for small samples (see Figure A2).
Table 1: Inverse Shape Estimates for Different Distributions
Mean Std.Dev. Q25 Median Q75
Normal–Normal $0.0416$ $0.0319$ 0.0104-0.0104 $0.0425$ $0.0954$
t(5)t(5)–Normal $0.1432$ $0.0363$ $0.0814$ $0.1431$ $0.2028$
Normal–$t(5)$ $0.1706$ $0.0362$ $0.1110$ $0.1697$ $0.2316$
$t(5)t(5)$ $0.2120$ $0.0384$ $0.1497$ $0.2112$ $0.2758$
$1,000$ repetitions for $N=100$ observations.
4.1.2 Location and Scale Estimation
We evaluate whether estimation of location and scale parameters with block maximum MLE can accurately capture the true distribution. (Results are provided in Figure A3.) The finite sample adjusted location parameter works well and shows convergence to the true value, while the scale exhibit slight downward bias that is consistent with known limitations of MLE for EVDs (Dombry & Ferreira, 2019). Panel (a) in Figure A3 summarises the effectiveness of the MLE in the Gumbel case, allowing us to conduct hypothesis tests.
4.2 Applications
We investigate several real-world datasets — two applications from economics and biology, as well as four well-known machine learning benchmarks.
4.2.1 Economic Development and Geography
We re-examine the controversial finding that rugged terrain benefits African economies when compared to the rest of the world (Nunn, 2020). Kuschnig et al. (2021) identify the Seychelles, coupled with any of Rwanda, Lesotho, Eswatini, and the Comoros, as influential, removing significance of the estimate of interest.
Our results decisively resolve this controversy. Table 2 reveals the Seychelles as excessively influential on $\hat{\theta}_{\text{rugged}}$, both individually ($p<0.001$) and in combination with other outliers except for Lesotho. This confirms the suspected confounding from the size of nations, lending statistical rigor to prior concerns.
Table 2: Influence of Ruggedness on log(GDP per capita in 2000)
Influential Set $\Delta({\mathbb{S}})$ $\hat{\tilde{a}}$ $\hat{b}$ $p$-value
Seychelles $0.077$ $0.020$ $0.004$ $1e^{-16}$
Seychelles + Lesotho $0.046$ $0.036$ $0.007$ $0.216$
Seychelles + Rwanda $0.070$ $0.028$ $0.006$ $0.001$
Seychelles + Eswatini $0.077$ $0.020$ $0.004$ $1e^{-16}$
Seychelles + Comoros $0.061$ $0.028$ $0.006$ $0.004$
4.2.2 Sparrow Morphology — Head and Beak Size
We analyze the relation between head and tarsus length in saltmarsh sparrows, based on measurements of $N=1295$ sparrows with known outliers (Gjerdrum et al., 2008; Zuur et al., 2010). The baseline regression yields $\hat{\theta}=0.011$ with a standard error of $(.030),$ implying a relation that is statistically indistinguishable from zero.
However, a curious data point moves the estimate to $0.219(.029)$, turning the estimate significantly positive. An additional data point further moves the estimate to $0.288(.032).$ These extreme impacts from a vanishing fraction of the sample are deemed excessive by our approach at any conventional significance level (both p<0.001p<0.001). [^3] A possible explanation for this excessive influence are data entry errors: The first observation (an outlier in both head and tarsus size) may have the two (adjacent) features mixed up — when swapped, they would fit well into overall averages. The second observation (an outlier in one feature) stands out with both values being equal up to the one significant digit.
4.2.3 Machine Learning Benchmarks
We apply our framework to four widely-used regression benchmarks: Law School, Adult Income, Boston Housing, and Communities & Crime. For each dataset, we identify a most influential set of interest and test for excessive influence.
  • Law School ($N=20,800$): We examine the coefficient for the ‘Other’ race indicator, with $378$ relevant samples. We consider two sets: $77$ data points that move the estimate from 0.0412(.0144)-0.0412\ (.0144) to 0.1117(.0159),0.1117\ (.0159), creating a significant estimate with flipped sign, and $17$ data points that reduce the estimate to 0.0223(.0097).-0.0223\ (.0097). Our approach indicates that the influence larger set’s influence falls within expected variation, while the smaller set exhibits statistically excessive influence ($p=0.019$).
  • Adult Income ($N=32,561$): We investigate the top 1% most influential sets ($325$ points) that shift the ‘Male’ indicator from $\hat{\theta}=0.062,$ either raising it to $0.0992$ or decreasing it to $0.0214.$ Despite these considerable shifts from a small fraction of the data, neither is deemed excessively influential by our approach.
  • Boston Housing ($N=506$): We focus on the effect of crime rate on house values. The baseline (highly significant) coefficient 0.1080(.0329)-0.1080\ (.0329) is rendered insignificant at 0.0352(.0556)-0.0352\ (.0556) after excluding just $6$ observations. In this case, the underlying EVD is Frechét with inverse shape $\xi^{-1}=0.29$ due to the heavy tail of the crime variable. The set’s influence is highly significant ($p=0.001$), indicating excessive influence.
  • Communities & Crime ($N=1,994$): We investigate $2$ and $2$ data points with substantial influence on the relation between race and crime rates. The complete set is not extreme, as the points cancel each other out. After exclusion, the first subset of two increases the coefficient by more than 22%, which is deemed excessive $p<0.001.$ When re-estimating after their exclusion, the second set decreases the estimate by more than 10% and is deemed excessive at the 5% level ($p=0.014.$ (See Table A2 for details.)
5 Discussion
Our analysis provides the first rigorous statistical framework for assessing when most influential sets represent genuine problems rather than natural sampling variation. By establishing that maximal influence follows predictable extreme value distributions, we enable practitioners to move beyond ad-hoc rules and domain-specific judgment. The key insight is that maximal influence fundamentally depends on the nature of the sets considered and the tail behavior of the underlying data.
This work addresses a critical gap in interpretable machine learning, where theoretical foundations for influential set analysis have been lacking. We show that maximal influence depends on the tail properties of the underlying data—when tails are heavy, influence patterns become unpredictable, requiring robust estimation methods or acceptance of the inherent instability. For distributions with moderate tail behavior, our results provide tight theoretical bounds.
Our approach integrates naturally with the literature on influential set selection (Broderick et al., 2021; Fisher et al., 2023; Hu et al., 2024), which has lacked conclusive theoretical guidance. While influential sets share connections with established diagnostics such as Cook’s distance and leverage scores, or methods such as robust regression (Huber & Ronchetti, 2009), they fill a unique role. By directly relating to quantities of interest, they allow practitioners to discover and analyze sets of data that deviate from dominant patterns and yield insights that other methods obscure (see, e.g., Kuschnig et al., 2021, for comparisons). Our results provide the theoretical foundations for analyzing these influential sets. [^4] We can also clarify the applicability of the common $2\sqrt{N}$ threshold for coefficient influence (Belsley et al., 1980). While imprecise for most influential observations — missing the selection procedure that necessitates extreme value theory — it proves asymptotically accurate for randomly selected observations.
5.1 Limitations and Future Work
Our analysis operates within linear regression — a foundational setting for theory and modern ML, but limited to contexts where interpretability is valued (Rudin, 2019; Roscher et al., 2020). While we focus on regression coefficients, an extension predictions follows trivially through $\hat{Y}=X\hat{\beta}$. Extending our results to generalized linear, tree-based, or non-parametric models, requires further developments.
The asymptotic arguments presented here provide important insights, but come with inherent limitations. Our analysis leverages independence between features and the residuals, which can be restrictive in practice where dependence affects influence patterns. While we address these concerns in controlled settings, the gap between theory and finite-sample behavior warrants investigation.
Several methodological improvements could strengthen practical performance. Estimation of extreme value parameters could be enhanced through setting-specific methods and improved bias correction (Dombry & Ferreira, 2019). The efficient selection of most influential sets themselves remains an open challenge (Hu et al., 2024; Huang et al., 2025) with direct implications for our procedure.
5.2 Broader Implications
This work enables more reliable and transparent decision-making across domains where linear models remain the method of choice. Principled tools for understanding data points that drive model behavior are crucial for building trustworthy systems. Applications span fairness assessments, where influential subsets can reveal algorithmic bias, to causal inference settings, including randomized controlled trials quasi-experimental econometric analyses where small data subset can fundamentally alter estimates.
Importantly, we reframe influence as a natural and informative feature of data that requires appropriate treatment rather than a probem to be fixed. Influential sets can represent genuine heterogeneity or important edge cases that should inform model development. This perspective emables more nuanced approaches to data processing and model validation, where more information is preserved and assessed through principled statistical inference rather than discarded based on rules of thumb.
6 Conclusion
We developed a statistical framework that transforms the assessment of most influential sets from art to science. By deriving the extreme value distributions of maximal influence, we enable rigorous hypothesis testing to distinguish excessive influence from natural variation. Applications across economics, biology, and machine learning benchmarks demonstrate the practical utility of our approach.
Our method offers clear guidance to practitioners — when small sets overturn results of interest, our tests reveal whether this influence is statistically excessive. This enables more robust and transparent decision-making in settings where reliability matters, from medical trials to policy evaluation to algorithmic systems. By providing theoretical foundations for influential set analysis, this work advances both the theory and practice of interpretable machine learning.
Reproducibility Statement
Theoretical results are elaborated upon in the Appendix, where proofs are elaborated upon via several Lemmata. Datasets can be obtained from the cited sources, and code that generates the results will be made available.
Statement on LLM Use
Large language models were used to (i) aid and polish writing, (ii) retrieve and discover related work, and (iii) check results for apparent mistakes.
Appendix A1 Exact Influence Formulas
Proposition 2.
Consider the influence in Equation 3. We can simplify this expression to
$\Delta\left(\{i\}\right)=\frac{x_{i}r_{i}}{\sum_{n\neq i}x_{n}^{2}}.$
Proof.
Let $A_{i}\mathrel{\raisebox{0.0pt}{:}}=x_{i}^{2}$ and $D\coloneqq\sum_{n}^{N}X_{n}^{2},$ such that $D_{-\{i\}}=\sum_{n\neq i}^{N}x_{n}^{2}.$ Then
$\Delta\left(\{i\}\right)=\frac{\frac{\sqrt{A_{i}}r_{i}}{(A_{i}+D_{-\{i\}})}}{1-\frac{A_{i}}{(A_{i}+D_{-\{i\}})}}$ $=\frac{(A_{i}^{1.5}\cdot r_{i}+A_{i}^{0.5}\cdot r_{i}\cdot D_{-\{i\}})}{D_{-\{i\}}\left(A_{i}+D_{-\{i\}}\right)}$
$=\frac{r_{i}\cdot x_{i}\cdot\left(A_{i}+D_{-\{i\}}\right)}{D_{-\{i\}}\cdot\left(A_{i}+D_{-\{i\}}\right)}=\frac{r_{i}\cdot x_{i}}{\sum_{n\neq i}^{N}x_{n}^{2}}.$
A1.1 Recursion
It helps to know the influence on the residual, leverage, and full hat matrix:
$(r_{i})_{-\{j\}}$ $=r_{i}+x_{i}\,\Delta\left(\{j\}\right),$
$(h_{i})_{-\{j\}}$ $=x_{i}^{2}/\sum_{n\neq j}x_{n}^{2},$
$(h_{ij})_{-\{k\}}$ $=h_{ij}+\frac{h_{ik}h_{kj}}{1-h_{k}}.$
Proposition 3.
Let ${\mathbb{S}}=\left\{1,\dots,K\right\}.$ Then we can recursively define $\Delta\left({\mathbb{S}}\right)$ as
$\Delta\left({\mathbb{S}}\right)$ $=\hat{\beta}-\hat{\beta}_{-{\mathbb{S}}}$
$=\Delta\left(\{1\}\right)+\Delta\left(\{2\}\right)_{-\{1\}}+\dots+\Delta\left(\{K\}\right)_{-{\mathbb{S}}\setminus K}.$
Proof.
Notice that
[…]
$=\Delta\left(\left\{j\right\}\right)+\Delta\left(\left\{i\right\}\right)_{-\{j\}}.$
Equivalent results trivially hold for larger sets. ∎
Now, we can provide the full details for the induction step in the proof of Proposition 1:
Proof.
$\Delta\left({\mathbb{S}}\right)=\frac{\sum_{k=1}^{K-1}x_{k}r_{k}}{D_{-{\mathbb{S}}}+x_{k}^{2}}+\frac{x_{k}r_{k}}{D_{-{\mathbb{S}}}}+\frac{x_{k}^{2}\sum_{k=1}^{K-1}x_{k}r_{k}}{D_{-{\mathbb{S}}}\left(D_{-{\mathbb{S}}}+x_{k}^{2}\right)}=\frac{\sum_{k=1}^{K}x_{k}r_{k}}{D_{-{\mathbb{S}}}},$
and
$\Delta\left({\mathbb{S}}\right)$ $=\frac{\sum_{k=1}^{K}x_{k}r_{k}}{D_{-{\mathbb{S}}}+x_{K+1}^{2}}+\frac{x_{K+1}r_{K+1}}{D_{-{\mathbb{S}}}}+\frac{x_{K+1}^{2}\sum_{k=1}^{K}x_{k}r_{k}}{D_{-{\mathbb{S}}}\left(D_{-{\mathbb{S}}}+x_{K+1}^{2}\right)}$
$=\frac{\sum_{i=k}^{K}x_{k}r_{k}\left(D_{-{\mathbb{S}}}+x_{K+1}^{2}\right)}{D_{-{\mathbb{S}}}\left(D_{-{\mathbb{S}}}+x_{K+1}^{2}\right)}+\frac{x_{K+1}r_{K+1}}{D_{-{\mathbb{S}}}}=\frac{\sum_{k=1}^{K+1}x_{k}r_{k}}{D_{-{\mathbb{S}}}}.$
Appendix A2 Lemma for the Inverse Sum of Squares
Lemma 1 (Asymptotic Normality of Inverse Sum of Squares).
Let $\{X_{i}\}_{i=1}^{\infty}$ be a sequence of independent and identically distributed (i.i.d.) random variables satisfying:
  1. 1.
    $\mathbb{E}[X_{1}^{4}]<\infty$ (finite fourth moment)
  2. 2.
    $\mathbb{E}[X_{1}^{2}]=\mu>0$ (positive second moment)
  3. 3.
    $\operatorname{Var}(X_{1}^{2})=\sigma^{2}>0$ (non-degenerate variance of squares)
Define $S_{n}=\sum_{i=1}^{n}X_{i}^{2}$ and $Y_{n}=S_{n}^{-1}$. Then $Y_{n}$ is asymptotically normal with:
$n^{3/2}\left(Y_{n}-\frac{1}{n\mu}\right)\xrightarrow{d}\mathcal{N}\left(0,\frac{\sigma^{2}}{\mu^{4}}\right)\quad\text{as }n\to\infty.$
Proof.
Define the sample mean of squares $\bar{X}_{n}^{(2)}=n^{-1}S_{n}$. By the Central Limit Theorem (CLT):
$\sqrt{n}\left(\bar{X}_{n}^{(2)}-\mu\right)\xrightarrow{d}\mathcal{N}(0,\sigma^{2}),$
where $\mu=\mathbb{E}[X_{1}^{2}]$ and $\sigma^{2}=\operatorname{Var}(X_{1}^{2})$ (finite by $\mathbb{E}[X_{1}^{4}]<\infty$).
Consider the transformation $g(x)=x^{-1}$, which is differentiable at $x=\mu>0$ with derivative $g^{\prime}(x)=-x^{-2}$. The Delta Method gives:
$\sqrt{n}\left(g(\bar{X}_{n}^{(2)})-g(\mu)\right)\xrightarrow{d}\mathcal{N}\left(0,\sigma^{2}\cdot[g^{\prime}(\mu)]^{2}\right).$
Substituting $g(\bar{X}_{n}^{(2)})=(\bar{X}_{n}^{(2)})^{-1}=n/S_{n}$ and $g(\mu)=\mu^{-1}$:
$\sqrt{n}\left(\frac{n}{S_{n}}-\frac{1}{\mu}\right)\xrightarrow{d}\mathcal{N}\left(0,\sigma^{2}\cdot(-\mu^{-2})^{2}\right)=\mathcal{N}\left(0,\frac{\sigma^{2}}{\mu^{4}}\right).$
Rewriting $n/S_{n}=nY_{n}$:
$\sqrt{n}\left(nY_{n}-\mu^{-1}\right)\xrightarrow{d}\mathcal{N}\left(0,\frac{\sigma^{2}}{\mu^{4}}\right).$
Factoring the left side:
$\sqrt{n}\left(nY_{n}-\mu^{-1}\right)$ $=n^{1/2}\cdot n\left(Y_{n}-\frac{1}{n\mu}\right)$
$=n^{3/2}\left(Y_{n}-\frac{1}{n\mu}\right).$
Thus:
$n^{3/2}\left(Y_{n}-\frac{1}{n\mu}\right)\xrightarrow{d}\mathcal{N}\left(0,\frac{\sigma^{2}}{\mu^{4}}\right).$
Appendix A3 Lemmata for the Product EVD
For notational simplicity let $S\coloneqq\sum_{i\in{\mathbb{S}}}X_{i}\cdot R_{i}$ and $T\coloneqq D^{-1}_{-{\mathbb{S}}},$ where $D=\sum_{n}X_{n}^{2}$. It holds for any realization $S=s\in\mathbb{R}$ and $T=t\in\mathbb{R^{+}}$. Further, let $\mathrm{MDA}(H)$ denote the maximum domain of attraction of an EVD $H$ where we write $Z\in\mathrm{MDA}(H)$. We specifically denote the Fréchet as $\Phi_{\alpha}$ and the Gumbel as $\Lambda.$ We are interested in the EVD of $\Delta=S\cdot T.$
A3.1 If $S\in\mathrm{MDA}(\Lambda)$
Lemma 2.
Let $T\in\mathrm{MDA}(\Lambda)$ and $S\in\mathrm{MDA}(\Phi_{a})$ with tail-coefficient $a>0$ and $S$ and $T$ being independent, then $\Delta\left({\mathbb{S}}\right)=S\cdot T\in\mathrm{MDA}(\Phi_{a})$ .
Proof.
Recall that for Gumbel tails ($S$) the survival function decays double-exponentially, i.e.,
$\mathbb{P}(S>s)\sim\exp\left(-\exp\left(\frac{s-\mu}{\beta}\right)\right)\quad\text{as }s\to\infty,$
while for the Fréchet tails ($T$) the survival function is regularly varying with index a-a, i.e.,
$\mathbb{P}(T>t)\sim t^{-a}L_{T}(t)\quad\text{as }t\to\infty,$
where $L_{T}(t)$ is a slowly varying function. The density satisfies:
$f_{T}(t)\sim at^{-a-1}L_{T}(t)\quad\text{as }t\to\infty.$
We are interested in the EVD of $\Delta$, i.e., $\mathbb{P}(\Delta>\delta)$. Since $\Delta=S\cdot T$ and $S$, $T$ are independent by construction:
$\mathbb{P}(\Delta>\delta)=\mathbb{P}(S\,T>\delta)=\int_{\mathbb{R}^{+}}\mathbb{P}(S>\delta/t)f_{T}(t)\,\mathrm{dt}.$
Next, we split the integral at $M>0$:
$\mathbb{P}(\Delta>\delta)=\underbrace{\int_{0}^{M}\mathbb{P}(S>\delta/t)f_{T}(t)\,\mathrm{dt}}_{I_{1}}+\underbrace{\int_{M}^{\infty}\mathbb{P}(S>\delta/t)f_{T}(t)\,\mathrm{dt}}_{I_{2}}.$
For fixed $M$, we have $I_{1}\to 0,$ as $\delta\to\infty$ since $\delta/t\to\infty$ and Gumbel tails decay faster than any polynomial, and the dominant term is $I_{2}$. Substitute $u=\delta/t$ ($t=\delta/u$, $\mathrm{dt}=-(\delta/u^{2})\,\mathrm{du}$), and we have
$I_{2}=\int_{M}^{\infty}\mathbb{P}(S>\delta/t)f_{T}(t)\,\mathrm{dt}=\int_{0}^{\delta/M}\mathbb{P}(S>u)f_{T}(\delta/u)\frac{\delta}{u^{2}}\,\mathrm{du}.$
Using the asymptotic form of $f_{T}$:
$f_{T}(\delta/u)\sim a(\delta/u)^{-a-1}L_{T}(\delta/u),$
we obtain
$I_{2}\sim\int_{0}^{\delta/M}\mathbb{P}(S>u)\left[a\left(\frac{\delta}{u}\right)^{-a-1}L_{T}\left(\frac{\delta}{u}\right)\right]\frac{\delta}{u^{2}}\,\mathrm{du}$
$=a\delta^{-a}\int_{0}^{\delta/M}\mathbb{P}(S>u)u^{a-1}L_{T}\left(\frac{\delta}{u}\right)\,\mathrm{du}.$
As $\delta\to\infty$, by Lemma 4 in Appendix A4, we obtain
$\int_{0}^{\delta/M}\mathbb{P}(S>u)u^{a-1}L_{T}\left(\frac{\delta}{u}\right)\,\mathrm{du}\sim L_{T}(\delta)\int_{0}^{\infty}\mathbb{P}(S>u)u^{a-1}\,\mathrm{du}.$ (A1)
The integral converges because:
  1. 1.
    near $u=0$ we have $\mathbb{P}(S>u)\approx 1$ and $u^{a-1}$ is integrable for $a>0$, and
  2. 2.
    as $u\to\infty$ the Gumbel decay dominates $u^{a-1}.$
Denote the constant
$C(a,S)=\int_{0}^{\infty}\mathbb{P}(S>u)u^{a-1}\,\mathrm{du}\in(0,\infty),$
then
$\mathbb{P}(\Delta>\delta)\sim a\delta^{-a}L_{T}(\delta)C(a,S)=\delta^{-a}\left(aC(a,S)L_{T}(\delta)\right).$
The term in parentheses is slowly varying in $\delta$ since $L_{T}(\delta)$ is slowly varying which concludes the proof. ∎
To summarize, the survival function $\mathbb{P}(\Delta>\delta)$ is regularly varying with index a,-a, and therefore, $\Delta$ has Fréchet tails with tail-coefficient $a.$
Corollary 2.
Following Lemma 2 and assuming a tail coefficient $a=\infty$ it follows that $S\sim\text{Gumbel}$ and thus $\Delta\left({\mathbb{S}}\right)=S\cdot T\in\mathrm{MDA}(\Lambda).$
Proof.
The result follows directly from properties of the Fréchet distribution. ∎
Lemma 3.
If $S\in\mathrm{MDA}(\Phi_{a})$ and $T\in\mathrm{MDA}(\Phi_{b})$ then $\Delta\left({\mathbb{S}}\right)\in\mathrm{MDA}(\Phi_{\min\{a,b\}})$.
Proof.
The proof of this follows directly from Lemma 1.3.1 on the convolution closure of distribution functions with regularly varying tails in Embrechts et al. (1997). ∎
Corollary 3 (Conditional EVD).
Further, if $S\in\mathrm{MDA}(E)$ for some EVD $E$, it holds that
$\Delta\left({\mathbb{S}}\right)\mid X_{-{\mathbb{S}}}\in\mathrm{MDA}(E),$
Appendix A4 Lemma for Asymptotic Equivalence
Lemma 4 (Asymptotic Equivalence Statement).
$\int_{0}^{\delta/M}\mathbb{P}(S>u)u^{a-1}L_{T}\left(\frac{\delta}{u}\right)\,\mathrm{du}\sim L_{T}(\delta)\int_{0}^{\infty}\mathbb{P}(S>u)u^{a-1}\,\mathrm{du},$
where $S$ has Gumbel tails, $L_{T}$ is slowly varying, $a>0$ is the tail coefficient, $M>0$ is a fixed constant.
Proof.
For clarity, we prove this result in five steps.
Step 1: Integral Splitting
Define
$I(\delta)=\int_{0}^{\delta/M}\mathbb{P}(S>u)u^{a-1}L_{T}\left(\frac{\delta}{u}\right)\,\mathrm{du}=I_{1}(\delta)+I_{2}(\delta),$
where
$I_{1}(\delta)$ $=\int_{0}^{1}\mathbb{P}(S>u)u^{a-1}L_{T}\left(\frac{\delta}{u}\right)\,\mathrm{du},$
$I_{2}(\delta)$ $=\int_{1}^{\delta/M}\mathbb{P}(S>u)u^{a-1}L_{T}\left(\frac{\delta}{u}\right)\,\mathrm{du}.$
Step 2: Analysis of $I_{1}(\delta)$ (Bounded Domain)
For $u\in(0,1],$ we have:
$\lim_{\delta\to\infty}\frac{I_{1}(\delta)}{L_{T}(\delta)}$ $=\lim_{\delta\to\infty}\int_{0}^{1}\mathbb{P}(S>u)u^{a-1}\frac{L_{T}(\delta/u)}{L_{T}(\delta)}\,\mathrm{du}$
$=\int_{0}^{1}\mathbb{P}(S>u)u^{a-1}\,\mathrm{du},$
by the Dominated Convergence Theorem (DCT):
  • Pointwise convergence: for fixed $u>0$, $\lim_{\delta\to\infty}\frac{L_{T}(\delta/u)}{L_{T}(\delta)}=1.$
  • Dominating function: by Potter’s theorem, for any $\delta>0$, there exists $C_{\delta}>0$ such that
    $\left|\frac{L_{T}(\delta/u)}{L_{T}(\delta)}\right|\leqslant C_{\delta}u^{-\delta}\quad\text{for all large }\delta.$
    Choose $\delta<a$ such that $u^{a-1-\delta}$ is integrable on $(0,1],$ then
    $\left|\mathbb{P}(S>u)u^{a-1}\frac{L_{T}(\delta/u)}{L_{T}(\delta)}\right|\leqslant C_{\delta}u^{a-1-\delta}\quad(\text{since }\mathbb{P}\leqslant 1),$
    and the dominating function $C_{\delta}u^{a-1-\delta}$ is integrable over $(0,1]$ for $a>\delta>0.$
Step 3: Analysis of $I_{2}(\delta)$ (Growing Domain)
For $u\in[1,\delta/M],$ we have
$\lim_{\delta\to\infty}\frac{I_{2}(\delta)}{L_{T}(\delta)}$ $=\lim_{\delta\to\infty}\int_{1}^{\delta/M}\mathbb{P}(S>u)u^{a-1}\frac{L_{T}(\delta/u)}{L_{T}(\delta)}\,\mathrm{du}$
$=\int_{1}^{\infty}\mathbb{P}(S>u)u^{a-1}\,\mathrm{du}\quad\text{ by the DCT.}$
  • Pointwise convergence: for fixed $u\geqslant 1$, $\lim_{\delta\to\infty}\frac{L_{T}(\delta/u)}{L_{T}(\delta)}=1$
  • Dominating function: by Potter’s theorem, for $\delta>0$:
    $\left|\frac{L_{T}(\delta/u)}{L_{T}(\delta)}\right|\leqslant C_{\delta}u^{\delta}\quad\text{for all large }\delta,u\geqslant 1.$
    Choose $\delta$ small enough such that $k=a-1+\delta>0,$ and
    $\int_{1}^{\infty}\mathbb{P}(S>u)u^{k}\,\mathrm{du}<\infty,$
    since the Gumbel decay dominates. Then
    $\left|\mathbb{P}(S>u)u^{a-1}\frac{L_{T}(\delta/u)}{L_{T}(\delta)}\right|\leqslant C_{\delta}\mathbb{P}(S>u)u^{k},$
    and the dominating function $C_{\delta}\mathbb{P}(S>u)u^{k}$ is integrable over $1,\infty).$
  • Tail control: as $\delta\to\infty$, the upper limit $\delta/M\to\infty$ and
    $\int_{\delta/M}^{\infty}C_{\delta}\mathbb{P}(S>u)u^{k}\,\mathrm{du}\to 0.$
Step 4: Negligibility of Omitted Tail
The tail beyond $\delta/M$ is negligible:
$R(\delta)=\int_{\delta/M}^{\infty}\mathbb{P}(S>u)u^{a-1}L_{T}\left(\frac{\delta}{u}\right)\,\mathrm{du}.$
  • For $u\geqslant\delta/M$, we have $\delta/u\leqslant M$ s.t. is bounded on compact sets: $L_{T}(\delta/u)\leqslant C_{M}.$
  • By the Gumbel tail properties, there exist a $\theta>0$ s.t. $\mathbb{P}(S>u)\leqslant e^{-u^{\theta}}$ for large $u.$ Thus
    $|R(\delta)|\leqslant C_{M}\int_{\delta/M}^{\infty}e^{-u^{\theta}}u^{a-1}\,\mathrm{du}=o(1)\quad\text{as }\delta\to\infty.$
  • Since $L_{T}(\delta)\to\infty$ or is slowly varying, $R(\delta)=o(L_{T}(\delta))$
Step 5: Final Combination
Combining all results, we have
$\frac{I(\delta)}{L_{T}(\delta)}=$ $\frac{I_{1}(\delta)+I_{2}(\delta)+R(\delta)}{L_{T}(\delta)}$
$=$ $\frac{I_{1}(\delta)}{L_{T}(\delta)}+\frac{I_{2}(\delta)}{L_{T}(\delta)}+o(1)$
$\implies$ $\int_{0}^{1}\mathbb{P}(S>u)u^{a-1}\,\mathrm{du}+\int_{1}^{\infty}\mathbb{P}(S>u)u^{a-1}\,\mathrm{du}$
$=$ $\int_{0}^{\infty}\mathbb{P}(S>u)u^{a-1}\,\mathrm{du}.$
Appendix A5 Estimation
Figure A1: Visualization of Table 1. Most notable is the regime change between thin and polynomial tails. While for the Frechét cases the average MLE is statistically insignificantly different form another, they are all individually different from the Gumbel case. While the finite sample $D^{-1}_{-{\mathbb{S}}}$ has thicker than Normal tails, acting as a regularizer on $\Delta^{\max}$ and thereby resulting in slight underestimation of the theoretically suggested $\xi^{-1}=0.20$.
Table A1: Inverse Shape Estimates for Different Distributions and Sample Sizes
NN Distribution Mean Std.Dev. Q25 Median Q75
100 Normal–Normal $0.0416$ $0.0319$ 0.0104-0.0104 $0.0425$ $0.0954$
t(5)t(5)–Normal $0.1432$ $0.0363$ $0.0814$ $0.1431$ $0.2028$
Normal–$t(5)$ $0.1706$ $0.0362$ $0.1110$ $0.1697$ $0.2316$
$t(5)t(5)$ $0.2120$ $0.0384$ $0.1497$ $0.2112$ $0.2758$
200 Normal–Normal $0.0267$ $0.0318$ 0.0249-0.0249 $0.0272$ $0.0786$
t(5)t(5)–Normal $0.1394$ $0.0361$ $0.0774$ $0.1400$ $0.1967$
Normal–$t(5)$ $0.1666$ $0.0353$ $0.1091$ $0.1670$ $0.2241$
$t(5)t(5)$ $0.2033$ $0.0374$ $0.1406$ $0.2049$ $0.2627$
500 Normal–Normal $0.0147$ $0.0316$ 0.0358-0.0358 $0.0159$ $0.0635$
t(5)t(5)–Normal $0.1509$ $0.0387$ $0.0854$ $0.1528$ $0.2133$
Normal–$t(5)$ $0.1692$ $0.0367$ $0.1076$ $0.1696$ $0.2316$
$t(5)t(5)$ $0.2088$ $0.0382$ $0.1437$ $0.2096$ $0.2685$
1000 Normal–Normal $0.0101$ $0.0333$ 0.0451-0.0451 $0.0100$ $0.0641$
t(5)t(5)–Normal $0.1596$ $0.0378$ $0.0941$ $0.1603$ $0.2215$
Normal–$t(5)$ $0.1726$ $0.0360$ $0.1137$ $0.1727$ $0.2333$
$t(5)t(5)$ $0.2101$ $0.0374$ $0.1495$ $0.2123$ $0.2689$
2000 Normal–Normal $0.0106$ $0.0327$ 0.0431-0.0431 $0.0114$ $0.0626$
t(5)t(5)–Normal $0.1731$ $0.0371$ $0.1113$ $0.1740$ $0.2323$
Normal–$t(5)$ $0.1759$ $0.0359$ $0.1185$ $0.1762$ $0.2387$
$t(5)t(5)$ $0.2150$ $0.0375$ $0.1521$ $0.2139$ $0.2763$
All results based on 1000 repetitions.
Figure A2: Convergence of cases to the limiting distribution, visualizing Table A1. Solid lines are estimated means from $1000$ repetitions while the shaded area is $\pm 1SD$. The Normal–Normal case as well as the $t(5)t(5)$ case converge quickly to the theoretical $\xi^{-1}=0$ and $\xi^{-1}=0.2$ respectively. The convergence rate of the mixed cases converge slower yet reasonably fast. In particular, the t(5)t(5)–Normal case is visibly requiring more observations to stabilize $D^{-1}_{\mathbb{S}}$. In general this finding supports the relevance and applicability of our limiting results to finite samples.
Figure A3: Simulation exercise for the performance of the simple MLE based on block maxima, correcting for block size. While the corrected location parameter $\hat{\tilde{a}}$ is close to unbiased, the scale parameter $\hat{b}$ suffers some downward bias using simple block maxima, which is in line with Dombry & Ferreira (2019). However, for practical purposes the block maxima is expected to be fitting reasonably well, as visible in panel (a).
Appendix A6 Auxiliary Results for Case Studies
As mentioned in the main text, Table A2 summarizes the results for testing the preselected set and its subsets for significant influence of the percent of black population on the violent crimes committed per population.
Table A2: Influence of % Black Population on Violent Crimes
Set Composition Set Size $\Delta({\mathbb{S}})$ $\hat{\tilde{a}}$ $\hat{b}$ $p$-value
Full Set 44 $0.0214$ $0.0076$ $0.0029$ $0.4914$
1st Partial 22 $0.0456$ $0.0050$ $0.0021$ $7.62e^{-7}$
2nd Partial after excl. 1st 22 0.0241-0.0241 $0.0051$ $0.0022$ $0.0141$

### v3

@@
 Abstract
-Small subsets of data with disproportionate influence on model outcomes can have dramatic impacts on conclusions, with a few data points sometimes overturning key findings.
-While recent work has developed methods to identify these most influential sets, no formal theory exists to determine when their influence reflects genuine problems rather than natural sampling variation.
-We address this gap by developing a principled framework for assessing the statistical significance of most influential sets.
-Our theoretical results characterize the extreme value distributions of maximal influence and enable rigorous hypothesis tests for excessive influence, replacing current ad-hoc sensitivity checks.
-We demonstrate the practical value of our approach through applications across economics, biology, and machine learning benchmarks.
+Small influential data subsets can dramatically impact model conclusions, with a few data points overturning key findings.
+While recent work identifies these most influential sets, there is no formal way to tell when maximum influence is excessive rather than expected under natural random sampling variation.
+We address this gap by developing a principled framework for most influential sets.
+Focusing on linear least-squares, we derive a convenient exact influence formula and identify the extreme value distributions of maximal influence — the heavy-tailed Fréchet for constant-size sets and heavy-tailed data, and the well-behaved Gumbel for growing sets or light tails.
+This allows us to conduct rigorous hypothesis tests for excessive influence.
+We demonstrate through applications across economics, biology, and machine learning benchmarks, resolving contested findings and replacing ad-hoc heuristics with rigorous inference.
 1 Introduction
 Machine learning (ML) models and statistical inferences can be highly sensitive to small subsets of data.
-In many applications, just a handful of samples can overturn key conclusions: two countries nullify the estimated effect of geography on development (Kuschnig et al., 2021), a single outlier flips the sign of a treatment effect (Broderick et al., 2021), or a small group of individuals drives disparate outcomes in algorithmic decision-making (Black & Fredrikson, 2021).
-These most influential sets — data subsets with the greatest influence on model predictions — are central to questions of interpretability, fairness, and robustness in modern machine learning (see, e.g., Black & Fredrikson, 2021; Chen et al., 2018; Chhabra et al., 2023; Ghorbani & Zou, 2019; Sattigeri et al., 2022).
+In many applications, just a handful of samples can overturn key conclusions: two countries nullify the estimated effect of geography on development (Kuschnig et al., 2021), a single outlier flips the sign of a treatment effect (Broderick et al., 2023), or a small group of individuals drives disparate outcomes in algorithmic decision-making (Black and Fredrikson, 2021).
+These most influential sets — data subsets with the greatest influence on model predictions — are central to questions of interprtability, fairness, and robustness in modern machine learning (see, e.g., Black and Fredrikson, 2021; Chen et al., 2018; Chhabra et al., 2023; Ghorbani and Zou, 2019; Sattigeri et al., 2022).
 Despite their practical importance, practitioners lack principled tools to assess whether a set’s influence is genuinely problematic.
-Current practice relies on domain expertise and ad-hoc sensitivity checks, while approximate methods such as influence functions (Koh & Liang, 2017; Fisher et al., 2023; Schioppa et al., 2023) systematically underestimate the impacts of sets and extreme cases (Basu et al., 2020; Koh et al., 2019).
-Recent work highlights both the promise and challenges of most influential subsets — small sets can drive results even in randomized trials (Broderick et al., 2021; Kuschnig et al., 2021), heuristic algorithms can fail in simple settings (Hu et al., 2024; Huang et al., 2025), and influence bounds remain an active area of research (Moitra & Rohatgi, 2023; Freund & Hopkins, 2023; Rubinstein & Hopkins, 2024).
+Current practice relies on heuristics, ad-hoc sensitivity checks, and domain expertise, while approximate methods such as influence functions (Koh and Liang, 2017; Fisher et al., 2023; Schioppa et al., 2023) systematically underestimate the impacts of sets and extreme cases (Basu et al., 2020; Koh et al., 2019).
+Recent work highlights both the promises and challenges of most influential subsets — small sets can drive results even in randomized trials (Broderick et al., 2023; Kuschnig et al., 2021), heuristic algorithms can fail in simple settings (Hu et al., 2024; Huang et al., 2025), and influence bounds remain an active area of research (Moitra and Rohatgi, 2023; Freund and Hopkins, 2023; Rubinstein and Hopkins, 2025).
 What remains missing is a principled method to distinguish natural sampling variation from genuinely excessive influence.
 We develop a statistical framework for assessing the significance of most influential sets.
-By focusing on linear regression — a tractable, interpretable, and widely-used setting that underlies many modern methods (Rudin, 2019) — we derive the exact asymptotic distributions of maximal influence.
+By focusing on linear regression — a tractable, interpretable, and widely-used setting that underlies many modern methods — we derive the exact asymptotic distributions of maximal influence.
 We show that two distinct regimes emerge depending on the size of the influential set: when the size is fixed, maximal influence converges to a heavy-tailed Fréchet distribution; when the size grows with the sample, maximal influence converges to a well-behaved Gumbel distribution.
@@
 We demonstrate their practical value via applications across economics, biology, and machine learning benchmarks, resolving ambiguous cases where influential sets drive contested findings.
-1.1 Contributions
-We present a comprehensive analysis of the influence of most influential sets, both theoretically and in practical applications. To summarize, our main contributions are:
+Contributions.
+We present a comprehensive analysis of the influence of most influential sets, both theoretically and in practical applications. Our main contributions are:
   1. 1.
@@
   2. 2.
-    Efficient implementation. We provide computationally efficient procedures for evaluating influence, making our approach practical for real-world applications.
+    Efficient implementation. We provide a closed-form formula for evaluating set influence, making our approach practical for real-world applications.
   3. 3.
     Empirical validation.
-We demonstrate the utility of our framework across domains, resolving the contested “Blessing of Bad Geography” in economics, assessing robustness in biological data of sparrow morphology, and auditing fairness in ML benchmark datasets.
-1.2 Outline
+We demonstrate the utility of our framework across domains, resolving the contested ‘Blessing of Bad Geography’ in economics, assessing robustness in biological data of sparrow morphology, and auditing fairness in ML benchmark datasets.
+To summarize — we provide the first rigorous theoretical results that allow us to interpret influence, and demonstrate practicality by resolving contested findings.
+Outline.
 The remainder of the paper is structured as follows.
@@
   • Scientific discovery: Rugged terrain generally hinders economic development, but not in Africa. What if this striking result is driven by just two small island nations?
-  • Fairness auditing: An algorithmic decision-making system produces vastly different outcomes for a protected group. What if this disparity is explained by only a handful of data points?
-  • Data cleaning: A single influential data point among a thousand samples flips a strong correlation to a null result. Should we trust the original finding or the one without the outlier?
+  • Fairness auditing: An algorithmic decision-making system produces different outcomes for a protected group. What if the disparity can be explained by only a handful of data points?
+  • Data cleaning: A single influential point among a thousand samples flips a strong correlation to a null result. Should we trust the original finding or the one without the outlier?
   • Data preprocessing: A microcredit experiment shows negligible outcome variations overall, except for a few outliers. How should we prepare and analyze the sample?
@@
 These sets are intuitive to interpret, directly tied to the quantity of interest, and provide a new dimension for assessing estimates by highlighting their support in the data.
+What has been unclear, however, is how to interpret and deal with influential sets.
+Existing influence methods quantify how much changes, but offer little guidance on interpretation.
+Current practice relies heavily on domain expertise and ad-hoc rules, lacking a statistically rigorous framework for judging influence.
+Heuristics, such as sign-flips or significance thresholds, can flag effects that are stable and miss genuinely problematic influence.
+We address this gap by quantifying whether observed maximum influence is statistically compatible with natural sampling variation.
 2.1 Formal Problem Statement
@@
 $\hat{\theta}=\underset{\theta\in\mathbb{R}^{Q}}{\arg\min}\sum_{n=1}^{N}\mathcal{L}\left(f(\theta,x_{n}),y_{n}\right).$
-Let $\left[N\right]=\left\{1,\dots,N\right\}$ and denote both an index set and its corresponding subsample as ${\mathbb{S}}\subset\left[N\right].$ For any subset $\mathbb{S}$}, we use a subscript $\hat{\theta}_{-{\mathbb{S}}}$ to denote a quantity $\theta$ without $\mathbb{S}$}, i.e.
+Let $\left[N\right]=\left\{1,\dots,N\right\}$ and denote both an index set and its corresponding subsample as ${\mathbb{S}}\subset\left[N\right].$ For any subset $\mathbb{S}$}, we use a subscript $\hat{\theta}_{-{\mathbb{S}}}$ to denote an estimate $\hat{\theta}$ without $\mathbb{S}$}, i.e.
 $\hat{\theta}_{-{\mathbb{S}}}=\underset{\theta\in\mathbb{R}^{Q}}{\arg\min}\sum_{n\not\in{\mathbb{S}}}\mathcal{L}\left(f(\theta,x_{n}),y_{n}\right).$
 Definition (Most Influential Set).
-For a positive integer $k\ll N,$ the $k$-most influential subset is
-${\mathbb{S}}_{k}^{\max}\coloneqq\underset{{\mathbb{S}}\subset\left[N\right],\lvert{\mathbb{S}}\rvert\leqslant k}{\arg\max}\Delta\left({\mathbb{S}};\phi\right),$
-where $\Delta\left({\mathbb{S}};\phi\right)=\phi(\hat{\theta})-\phi(\hat{\theta}_{-{\mathbb{S}}})$ is the influence of subset $\mathbb{S}$} on target function $\phi:\mathbb{R}^{Q}\mapsto\mathbb{R}.$ The maximum influence is denoted as $\Delta^{\max}=\Delta\left({\mathbb{S}}_{k}^{\max};\phi\right).$
+For a positive integer $k\ll N,$ the $k$-most influential set is
+${\mathbb{S}}_{k}^{\max}\coloneqq\underset{{\mathbb{S}}\subset\left[N\right],\lvert{\mathbb{S}}\rvert=k}{\arg\max}\Delta\left({\mathbb{S}};\phi\right),$
+where $\Delta\left({\mathbb{S}};\phi\right)=\phi(\hat{\theta})-\phi(\hat{\theta}_{-{\mathbb{S}}})$ is the influence of subset $\mathbb{S}$} on the scalar target function $\phi:\mathbb{R}^{Q}\mapsto\mathbb{R}.$ We denote the maximum influence as $\Delta^{\max}=\Delta\left({\mathbb{S}}_{k}^{\max};\phi\right).$
 Research Question.
-What is the probability distribution of $\Delta^{\max},$ and how can we use it to distinguish excessive influence from natural sampling variation?
+What is the probability distribution of $\Delta^{\max},$ and how can we distinguish excessive influence from natural sampling variation?
 2.2 Influence Functions vs. Exact Influence
-A common approach to study influence is via influence functions (Fisher et al., 2023). These are motivated by reweighing via the perturbation
-$\hat{\theta}(\epsilon;{\mathbb{S}})\coloneqq\underset{\theta\in\mathbb{R}^{Q}}{\arg\min}\frac{1}{N}\sum_{n\not\in{\mathbb{S}}}\mathcal{L}\left(f(\theta,x_{n}),y_{n}\right)+\epsilon\sum_{i\in{\mathbb{S}}}\mathcal{L}\left(f(\theta,x_{i}),y_{i}\right).$
-Setting $\epsilon=0$ recovers $\hat{\theta}$, while $\epsilon=-N^{-1}$ yields $\hat{\theta}_{-{\mathbb{S}}}.$ The influence function is the linear approximation at $\epsilon=0$:
-$\hat{\theta}(\epsilon;{\mathbb{S}})\approx\mathcal{I}({\mathbb{S}})\coloneqq\frac{\mathrm{d}\hat{\theta}(\epsilon;{\mathbb{S}})}{\mathrm{d}\epsilon}\bigg|_{\epsilon=0}.$ (1)
-This yields the following first-order estimate of influence:
-$\Delta\left({\mathbb{S}};\phi\right)\approx-N^{-1}\frac{\mathrm{d}\phi\left(\hat{\theta}(\epsilon;{\mathbb{S}})\right)}{\mathrm{d}\epsilon}\bigg|_{\epsilon=0}.$
+A common and related approach to study influence is via influence functions (Fisher et al., 2023; Hu et al., 2024; Koh and Liang, 2017). These are motivated by reweighing via the perturbation
+$\hat{\theta}(\epsilon;{\mathbb{S}})\coloneqq\underset{\theta\in\mathbb{R}^{Q}}{\arg\min}\frac{1}{N}\sum_{n=1}^{N}\mathcal{L}\left(f(\theta,x_{n}),y_{n}\right)+\epsilon\sum_{i\in{\mathbb{S}}}\mathcal{L}\left(f(\theta,x_{i}),y_{i}\right).$
+Setting $\epsilon=0$ recovers $\hat{\theta}$, while $\epsilon=-N^{-1}$ yields $\hat{\theta}_{-{\mathbb{S}}}.$ The influence function is the first-order linear approximation at $\epsilon=0.$
 While influence functions are computationally convenient, they are unreliable even for simple models (Basu et al., 2020; Hu et al., 2024; Huang et al., 2025; Koh et al., 2019).
@@
 This occurs because the first-order approximation cannot reflect higher-order effects from the interplay between data points or differential leverage scores.
-Exact Influence
-In this work, we focus on the exact influence of subsets in a tractable but ubiquitous setting.
-We avoid linear approximations to accurately assess the most influential sets of interest, where extreme behavior dominates and first-order approximations fail most dramatically.
+Exact Maximum Influence.
+We therefore derive an exact influence formula (in Section 3) and characterize the behavior of maximum influence (Section 3.1) to enable principled testing (Section 3.3).
+We focus on the tractable but ubiquitous linear setting, allowing us to accurately portray most influential sets, for which extreme behavior dominates and first-order approximations fail most dramatically.
+2.2.1 Extreme Value Theory
+Our goal is to characterize the behavior of $\Delta^{\max},$ the influence of the most influential set.
+This quantity is defined through maximization over all possible subsets, and its distribution is governed by extreme value theory rather than classical asymptotics.
+See Figure 1 for an illustration.
+When taking maxima over random quantities, three possible limiting distributions can emerge (Fisher and Tippett, 1928; Gnedenko, 1943): the well-behaved Gumbel (Type I), the heavy-tailed Fréchet (Type II), and the bounded Weibull (Type III).
+Figure 1: Illustration of the distribution of $\Delta(\{i\})$ for general observations, and the maximum influence $\Delta^{\max}$ over all possible $i,$ for $N\in\{25,50,100\}.$ One can clearly discern an upward shift and a substantial increase of the density in the tails.
+We need to determine which extreme value distribution (EVD) attracts $\Delta^{\max}.$ Specifically, we distinguish between the Gumbel distribution with exponential tails, and the Fréchet distribution with polynomial tails allowing for potentially arbitrarily large influence.
+(The Weibull distribution can be ruled out since influence is unbounded.) Once we establish the asymptotic distribution, we show how it applies to finite samples.
+2.3 Setting
+Consider the standard linear regression model, where $f$ is a linear function relating random features $X$ to the outcome $Y$ via the parameter vector $\theta$.
+The ordinary least-squares (OLS) population estimator of $\theta$ is
+$\tilde{\theta}\coloneqq\mathbb{E}[XX^{\prime}]^{-1}\mathbb{E}[XY]$
+yielding fitted values $\hat{Y}=X^{\prime}\tilde{\theta}$ and the associated residuals $R=Y-X^{\prime}\tilde{\theta}$.
+Stacking $N$ observed training samples yields the design matrix $\mathbf{X}\in\mathbb{R}^{N\times P}$ and outcome vector $\mathbf{y}\in\mathbb{R}^{N}.$ For OLS, we assume that $\mathbf{X}^{\prime}\mathbf{X}$ is invertible and remains so after removing any subset.
+[^1] This is not necessary for our results, as they extend to ridge regression, where the penalization parameter $\lambda>0$ in $\mathbf{X}^{\prime}\mathbf{X}+\lambda\mathbf{I}$ creates a ridge that guarantees invertibility.
+The sample OLS estimator is then
+$\hat{\theta}\coloneqq\underset{\theta}{\arg\min}\lVert\mathbf{y}-\mathbf{X}\theta\rVert^{2}=\left(\mathbf{X}^{\prime}\mathbf{X}\right)^{-1}\mathbf{X}^{\prime}\mathbf{y},$
+yielding predictions $\hat{\mathbf{y}}=\mathbf{X}\hat{\theta}$ and residuals $\mathbf{r}=\mathbf{y}-\hat{\mathbf{y}}$.
+For our main results, we consider least-squares estimates of the regression coefficients, with and without penalization.
+For illustration, however, we will consider a univariate model with one positive coefficient of interest and target function $\phi(\theta)=\theta_{1}.$ This eases notation considerably and comes without loss of generality — theoretical results apply to more general settings.
+[^2] Multiple features can be factored out via the Frisch-Waugh-Lovell theorem under mild assumptions, leaving an equivalent univariate regression.
+A different sign can be accommodated by simply sign-flipping the feature of interest or adjusting the target function $\phi$ accordingly.
 3 Proposed Approach
-Consider the standard linear regression model, where $f$ is a linear function relating features $X\in\mathbb{R}^{P}$ to the outcome $Y\in\mathbb{R}.$ Stacking the observed training sample yields the design matrix $\mathbf{X}\in\mathbb{R}^{N\times P}$ and outcome vector $\mathbf{y}\in\mathbb{R}^{N}.$ We assume that $\mathbf{X}^{\prime}\mathbf{X}$ is invertible and remains so after removing any subset.
-[^1] Alternatively, one may consider the penalized estimator with $\mathbf{X}^{\prime}\mathbf{X}+\lambda\mathbf{I},$ where $\lambda>0$ creates a ridge that guarantees invertibility.
-Let $\mathbf{H}=\mathbf{X}\left(\mathbf{X}^{\prime}\mathbf{X}\right)^{-1}\mathbf{X}^{\prime}$ denote the hat matrix, with leverage scores along its diagonal $\mathbf{h}$.
-The ordinary least squares (OLS) estimator is
-$\hat{\theta}=\underset{\theta}{\arg\min}\lVert Y-X\theta\rVert^{2}=\left(\mathbf{X}^{\prime}\mathbf{X}\right)^{-1}\mathbf{X}^{\prime}\mathbf{y},$
-with predictions $\hat{\mathbf{y}}=\mathbf{X}\hat{\theta}$ and residuals $\mathbf{r}=\mathbf{y}-\hat{\mathbf{y}}$.
-We focus on the influence of subsets on a particular coefficient of interest for interpretation.
-For simplicity, we assume a univariate model with a positive coefficient, orthogonalize features where necessary, and set the target function to $\phi(\theta)=\theta_{1}.$
-3.1 Exact Influence Formulas
-The influence of a single observation $i$ is well-known (Belsley et al., 1980; Cook, 1979) to be
-$\Delta\left(\{i\}\right)=\frac{x_{i}r_{i}}{\sum_{n}x_{n}^{2}}\frac{1}{1-h_{i}},$ (2)
+The influence of a single observation $i$ is well-known (Belsley et al., 1980; Cook, 1979; Walker and Birch, 1988) to be
+$\Delta\left(\{i\}\right)=\frac{\left(\mathbf{X}^{\prime}\mathbf{X}\right)^{-1}\mathbf{x}_{i}r_{i}}{1-h_{i}},$
 where $h_{i}$ is the leverage score and $r_{i}$ the residual of observation $i.$
-For sets of observations, we can derive a closed-form expression via recursive application of Equation 2 and updating formulae for $r_{i}$ and $h_{i}.$ (Details are provided in the Appendix.) The following closed form solution is particularly convenient:
+We extend this result to be (i) applicable to sets of observations, and (ii) analytically convenient.
 Proposition 1.
-The influence of set $\mathbb{S}$} on $\hat{\theta}$ is
-$\Delta\left({\mathbb{S}}\right)=\frac{\sum_{i\in{\mathbb{S}}}x_{i}r_{i}}{\sum_{n\not\in{\mathbb{S}}}x_{n}^{2}}.$ (3)
-Proof sketch.
-For a single observation, one can show that $\Delta\left(i\right)=\frac{x_{i}r_{i}}{\sum_{n\neq i}x_{n}^{2}}.$ Then, let ${\mathbb{S}}=\left\{1,2\right\}$ and define $D\coloneqq\sum_{n}x_{n}^{2}$ for simplicity.
-Evidently,
-$\Delta\left(\{1,2\}\right)$ $=\frac{x_{1}r_{1}}{D_{-\{1,2\}}+x_{2}^{2}}+\frac{x_{2}\left(r_{2}+x_{2}\Delta\left(\{1\}\right)\right)}{D_{-\{1,2\}}}$
-$=\frac{x_{1}r_{1}}{D_{-\{1,2\}}+x_{2}^{2}}+\frac{x_{2}r_{2}}{D_{-\{1,2\}}}+\frac{x_{2}^{2}\left(x_{1}r_{1}\right)}{D_{-\{1,2\}}\left(D_{-\{1,2\}}+x_{2}^{2}\right)}$
-$=\frac{\left(D_{-\{1,2\}}+x_{2}^{2}\right)x_{1}r_{1}}{D_{-\{1,2\}}\left(D_{-\{1,2\}}+x_{2}^{2}\right)}+\frac{x_{2}r_{2}}{D_{-\{1,2\}}}=\frac{x_{1}r_{1}+x_{2}r_{2}}{D_{-\{1,2\}}},$
-where the second term in the first line corrects $r_{2}$ to reflect the removal of observation $1,$ which the second line expands.
-The third line merges terms one and three by transforming to a common denominator, and the fourth line simplifies the expression.
-Assuming this identity holds for $\lvert{\mathbb{S}}\rvert=K$ we can show by induction that it holds for $\lvert{\mathbb{S}}\rvert=K+1,$ and the result follows.
-Details are provided in the Appendix.
-∎
-Proposition 1 elegantly reveals the additive structure of individual contributions in the numerator and the multiplicative adjustment from the remaining data in the denominator.
-This representation enables efficient computation without explicitly forming leverage scores for each subset, making our approach computationally tractable for large datasets.
-3.2 Extreme value distribution
-We now turn to the distribution of $\Delta({\mathbb{S}})$ for the most influential set, ${\mathbb{S}}_{k}^{\max}.$ Since this quantity is defined by an extremal operation (maximization over all possible subsets), its asymptotic behavior is governed by extreme value theory.
-Specifically, we seek the limiting extreme value distribution (EVD) $H$ such that $\Delta^{\max}\in\mathrm{MDA}\left(H\right),$ i.e., $\Delta^{\max}$ lies in the maximum domain of attraction of $H.$
+The influence of some set $\mathbb{S}$} on the least-squares estimator $\hat{\theta}$ is
+$\Delta\left({\mathbb{S}}\right)=\left(\mathbf{X}_{-{\mathbb{S}}}^{\prime}\mathbf{X}_{-{\mathbb{S}}}+\lambda\mathbf{I}_{P}\right)^{-1}\mathbf{X}_{{\mathbb{S}}}^{\prime}\mathbf{r}_{{\mathbb{S}}},$ (1)
+where $\lambda\geqslant 0$ is an optional penalization parameter.
+Proof.
+Partition the full normal equations into $\mathbb{S}$} and $\mathbb{S}$}, and subtract $\mathbf{X}_{\mathbb{S}}^{\prime}\mathbf{X}_{\mathbb{S}}\hat{\theta}$:
+$(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})\,\hat{\theta}$ $=\mathbf{X}^{\prime}\mathbf{y}$
+$(\mathbf{X}_{-{\mathbb{S}}}^{\prime}\mathbf{X}_{-{\mathbb{S}}}+\mathbf{X}_{\mathbb{S}}^{\prime}\mathbf{X}_{\mathbb{S}}+\lambda\mathbf{I})\,\hat{\theta}$ $=\mathbf{X}_{-{\mathbb{S}}}^{\prime}\mathbf{y}_{-{\mathbb{S}}}+\mathbf{X}_{{\mathbb{S}}}^{\prime}\mathbf{y}_{\mathbb{S}}$
+[…]
+The key element stems from $\mathbf{r}_{\mathbb{S}}=\mathbf{y}_{\mathbb{S}}-\mathbf{X}_{\mathbb{S}}\,\hat{\theta}$, and the result follows from subtracting the leave-out normal equations and inverting:
+$(\mathbf{X}_{-{\mathbb{S}}}^{\prime}\mathbf{X}_{-{\mathbb{S}}}+\lambda\mathbf{I})\,\hat{\theta}$ $=\mathbf{X}_{-{\mathbb{S}}}^{\prime}\mathbf{y}_{-{\mathbb{S}}}+\mathbf{X}_{\mathbb{S}}^{\prime}\mathbf{r}_{\mathbb{S}}$
+$(\mathbf{X}_{-{\mathbb{S}}}^{\prime}\mathbf{X}_{-{\mathbb{S}}}+\lambda\mathbf{I})\,\left(\hat{\theta}-\hat{\theta}_{-{\mathbb{S}}}\right)$ $=\mathbf{X}_{\mathbb{S}}^{\prime}\mathbf{r}_{\mathbb{S}}$
+$\Delta({\mathbb{S}})$ $=(\mathbf{X}_{-{\mathbb{S}}}^{\prime}\mathbf{X}_{-{\mathbb{S}}}+\lambda\mathbf{I})^{-1}\mathbf{X}_{\mathbb{S}}^{\prime}\mathbf{r}_{\mathbb{S}}.$
+We thank an anonymous reviewer for pointing us towards this elegant proof. ∎
+Proposition 1 reveals the additive structure of individual contributions in the numerator, as well as the multiplicative adjustment from the denominator.
+It provides an exact closed-form expression that avoids re-fitting the model for each candidate subset.
+3.1 Extreme Value Distributions
+We now turn to finding the distribution of $\Delta({\mathbb{S}})$ for the most influential set, ${\mathbb{S}}_{k}^{\max}.$ Since this quantity is defined by an extremal operation (maximization over all possible subsets), its asymptotic behavior is governed by extreme value theory.
+Specifically, we seek the limiting EVD $H$ such that $\Delta^{\max}\in\mathrm{MDA}\left(H\right),$ i.e., $\Delta^{\max}$ lies in the maximum domain of attraction of $H.$
 Two canonical EVDs are of particular interest: the Fréchet (Type II) distribution $\Phi_{\alpha}$ for heavy-tailed variables, and the Gumbel (Type I) distribution $\Lambda$ for light-tailed variables.
@@
   2. 2.
-    Relative-size sets: $k$ grows proportionally with $N,$ i.e., $k=pN$ for some $p\in(0,1).$
-Both regimes appear in practical applications (see, e.g, Broderick et al., 2021; Kuschnig et al., 2021), but they yield fundamentally different asymptotic behavior with important implications for significance testing.
+    Growing-size sets: $k$ grows as $N\to\infty$ and $k/N\to 0$
+Both regimes have been considered in practical applications (see, e.g, Broderick et al., 2023; Kuschnig et al., 2021), and — as we will show next — they yield fundamentally different asymptotic behavior with important implications for the interpretation of influence.
 3.1.1 Constant-size Sets
 Theorem 1 (EVD for constant-size sets).
-Suppose $\mathbb{E}\left[X^{2}\right]<\infty,$ and that $X_{i},R_{i}$ have polynomial tails with coefficients $\xi_{x},\xi_{r}<\infty$.
-If $\lvert{\mathbb{S}}_{k}^{\max}\rvert$ remains constant as $N\to\infty,$ then
+Suppose $\mathbb{E}\left[X^{2}\right]<\infty,$ and that the heavier tail of $X_{i},R_{i}$ decays at polynomial speed with coefficients $\xi_{x},\xi_{r}$ where $\min\{\xi_{x},\xi_{r}\}<\infty.$ If $\lvert{\mathbb{S}}_{k}^{\max}\rvert$ remains constant as $N\to\infty,$ then
 $\lim_{N\to\infty}\Delta^{\max}\sim\text{Fréchet}(a,b,\xi),$
@@
 Let $C\coloneqq\sum_{i\in{\mathbb{S}}}X_{i}R_{i}$ and $D\coloneqq\sum_{n=1}^{N}X_{n}^{2}.$ Notice that $C$ and $D^{-1}_{-{\mathbb{S}}}$ are asymptotically independent.
-Since $X_{i}$ and $R_{i}$ have polynomial tails with coefficients $\xi_{x},\xi_{r},$ their product satisfies $C\in\mathrm{MDA}(\Phi_{\xi}),$ with $\xi=\min\{\xi_{x},\xi_{r}\},$ and its upper tail behaves like the tail of $\max\{X_{i}R_{i}\}$ for $i\in{\mathbb{S}}_{k}^{\max}.$ Lemma 1 shows that the inverse sum $D^{-1}_{-{\mathbb{S}}}\in\mathrm{MDA}(\Lambda)$, and the product $CD^{-1}_{-{\mathbb{S}}}$ inherits the Fréchet behavior from $C$ by Lemma 2.
-See the Appendix for details.
+Since $X_{i}$ and $R_{i}$ have polynomial tails with coefficients $\xi_{x},\xi_{r},$ their product $C\in\mathrm{MDA}(\Phi_{\xi}),$ with $\xi=\min\{\xi_{x},\xi_{r}\},$ since its upper tail behaves like the tail of $\max\{X_{i}R_{i}\}$ for $i\in{\mathbb{S}}_{k}^{\max}.$ Lemma 1 shows that the inverse sum $D^{-1}_{-{\mathbb{S}}}\in\mathrm{MDA}(\Lambda)$, and the product $CD^{-1}_{-{\mathbb{S}}}$ inherits the Fréchet behavior from $C$ by Lemma 2.
 ∎
-This result shows that for constant-size sets, $\Delta^{\max}$ exhibits heavy-tailed Fréchet behavior, implying that even a few observations can exert extreme influence with non-negligible probability.
+This result shows that, for constant-size sets, $\Delta^{\max}$ inherits tail behavior from the heavier tail of $R$ and $X.$ If one of them is sufficiently heavy-tailed, even small sets can exert extreme influence with non-negligible probability.
+Corollary 1 simplifies Theorem 1 in absence of heavy tails.
 Corollary 1.
-If the tail coefficients of both $X_{i}$ and $R_{i}$ is infinite, then
+If the tail coefficients of both $X_{i}$ and $R_{i}$ are infinite, then
 $\lim_{N\to\infty}\Delta^{\max}\sim\text{Gumbel}(a,b).$
-3.2.2 Relative-size Sets
-When the most influential set grows proportionally with the sample size, the central limit theorem dominates the asymptotic behavior:
-Theorem 2 (EVD for relative-size sets).
-If $\{X_{n}R_{n}\}_{n=1}^{N}$ satisfies the conditions of a central limit theorem (CLT) and $\lvert{\mathbb{S}}_{k}^{\max}\rvert$ grows proportionally with $N,$ then
+3.2 Growing Sets
+When the most influential set grows ‘slow enough’ with the sample size, the central limit theorem (CLT) dominates the asymptotic behavior:
+Theorem 2 (EVD for growing sets).
+If $\{X_{n}R_{n}\}_{n=1}^{N}$ satisfies the conditions of a CLT and $\lvert{\mathbb{S}}_{k}^{\max}\rvert$ grows at $o(N)$ but faster than $O(1),$ then
 $\lim_{N\to\infty}\Delta^{\max}\sim\text{Gumbel}(a,b).$
 Proof sketch.
-When $\lvert{\mathbb{S}}_{k}^{\max}\rvert=pN,$ for $p\ll 1$, the numerator $C$ grows at the rate $\mathcal{O}(N).$ By the CLT, $C/\sqrt{N}\sim\mathcal{N}(\mu,\sigma^{2})$ as $N\to\infty$.
-Hence, the product $CD^{-1}_{-{\mathbb{S}}}$ lies in the maximum domain of attraction of the Gumbel distribution, following Lemma 2 and Corollary 2.
-See the Appendix for details.
+Let $m_{k}=\lvert{\mathbb{S}}_{k}^{\max}\rvert$.
+By assumption, $m_{k}\to\infty$ and $m_{k}=o(N)$ as $N\to\infty$.
+The numerator, $C,$ can be written as a partial sum of $m_{k}$ terms and grows at rate $\mathcal{O}(m_{k})$.
+Since $\{X_{n}R_{n}\}_{n=1}^{N}$ satisfies the conditions of a CLT, we obtain $\left(C-\mathbb{E}[C]\right)/\sqrt{m_{k}}\;\overset{d}{\longrightarrow}\;\mathcal{N}(\mu,\sigma^{2})$ as $N\to\infty$.
+Hence, following Lemma 2 and Corollary 2, the product $CD^{-1}_{-{\mathbb{S}}}$ lies in the maximum domain of attraction of the Gumbel distribution.
 ∎
-For relative-size sets, $\Delta^{\max}$ converges to a well-behaved Gumbel distribution with exponentially decaying tails, in contrast to the heavy-tailed Fréchet behavior of constant-size sets.
+This reveals a fundamental distinction: constant-size sets are dominated by the heaviest tail, while, for growing sets, $\Delta^{\max}$ converges to a well-behaved Gumbel distribution with exponentially decaying tails.
 This result holds regardless of the underlying distributions of $X$ and $R$ as long as the variance of $X_{i}\cdot R_{i}$ is finite.
-3.3 Implementation and Computation
-With theoretical results established, we turn towards practical implementation. Our procedure follows three steps:
+3.3 Implementation
+With the theoretical results established, we turn towards practical implementation.
+Assuming there is a most influential set of interest, [^3] This set can be obtained with any of the methods in the literature (Broderick et al., 2023; Kuschnig et al., 2021; Freund and Hopkins, 2023), and could, e.g., be the smallest set that achieves a sign-flip — a commonly considered heuristic cutoff.
+If no such heuristic is used and multiple tests are considered (e.g., for different coefficients or sets sizes), a multiple testing correction should be applied.
+Note that EVT controls the implicit search for the most influential set over the $\binom{N}{k}$ possible subsets.
+our procedure follows three steps:
   1. 1.
-    Determine the relevant extreme value distribution.
-We select between the Gumbel and Fréchet families based on the hypothesized set size and the tail behavior of $X$ and $R$.
-We estimate tail coefficients using maximum likelihood estimation (MLE; Smith, 1985; Bücher & Segers, 2017).
-If $1/\max\{\xi_{x},\xi_{r}\}$ is sufficiently close to zero, we default to the Gumbel distribution (per Corollary 1 and Theorem 2).
-Otherwise, we use the Fréchet distribution with shape parameter $\xi=\max\{\xi_{x},\xi_{r}\},$ following Theorem 1.222Small values of $\xi$ correspond to extremely heavy-tailed distributions where the variance ($\xi\leqslant 2$) or even the mean ($\xi\leqslant 1$) become infinite.
-Such extreme cases pose practical challenges for statistical inference.
+    Choose the EVD family.
+Our theoretical results guide the decision between the Gumbel and Fréchet families, which based on the hypothesized set size and the tail behavior of $X$ and $R$.
+For the latter, we can estimate tail coefficients using maximum likelihood estimation (MLE; Smith, 1985; Bücher and Segers, 2017).
+If $1/\xi$ is sufficiently close to zero, we default to the Gumbel distribution (per Corollary 1 and Theorem 2).
+Otherwise, we use the Fréchet distribution with shape parameter $\xi,$ following Theorem 1.444Small values of $\xi$ correspond to extremely heavy-tailed distributions where the variance ($\xi\leqslant 2$) or even the mean ($\xi\leqslant 1$) become infinite.
+Such cases pose practical challenges for statistical inference, and imply that arbitrarily large influence is possible in our case.
   2. 2.
-    Estimate location and scale parameters.
-We estimate the location and scale parameters $a,b$ using the block maxima method (Coles, 2001; De Haan & Ferreira, 2006).
-We divide the sample (excluding ${\mathbb{S}}_{k}^{\max}$ for robustness) into $M$ blocks of size $N/M$, compute $\Delta^{\max}$ for each block, and use MLE based on these draws.
-    Since selecting the maximum out of $N/M$ observations reduces the expected maximum compared to the full sample, in the Gumbel case a bias correction can be applied. More specifically, we know that
+    Estimate EVD parameters.
+With the EVD known, we estimate its location and scale parameters $a,b$ e.g., using the block maxima method (Coles, 2002; de Haan and Ferreira, 2006).
+For this, we divide the sample (excluding ${\mathbb{S}}_{k}^{\max}$ for robustness) into $M$ blocks of size $N/M$, compute $\Delta^{\max}$ for each block, and use MLE based on these draws.
+Since selecting the maximum out of $N/M$ observations reduces the expected maximum compared to the full sample, a bias correction can be applied for the Gumbel distribution, leveraging that the densities for sizes $N$ and $N/M$ are related by
     $F^{N}(x)\xrightarrow{d}\text{Gumbel}(a,b)\quad\text{and}\quad[F^{N/M}(x)]^{M}\xrightarrow{d}\text{Gumbel}(a,b),$
@@
   3. 3.
-    Perform hypothesis testing.
-We test the null hypothesis $H_{0}$ that the observed influence reflects natural sampling variation against the $H_{1}$ of excessive influence.
-Using the estimated parameters, we compute the $p$-value as $P(\Delta^{\max}\geqslant\delta_{\text{obs}})$ where $\delta_{\text{obs}}$ is the observed maximum influence.
-Computational Efficiency.
-Our procedure is computationally convenient, allowing for application to large and varied datasets.
-The maximum likelihood steps are simple and well-behaved, optimizing over only two parameters in the Gumbel case.
-The primary computational constraint stems from finding most influential sets — we need to approximate $\Delta^{\max}$ for the $M$ block maxima estimates (Price et al., 2022).
-For computational tractability, we use an adaptive greedy algorithm (Hu et al., 2024; Kuschnig et al., 2021) with complexity $\mathcal{O}(Mk)$ and considerably reduced runtime from our closed-form influence formula for sets in Proposition 1.
+    Perform hypothesis test.
+Finally, we test the null hypothesis $H_{0}$ that the observed influence reflects natural sampling variation against the alternative $H_{1}$ of excessive influence.
+Based on the estimated parameters, we can simply compute the $p$-value as $P(\Delta^{\max}\geqslant\delta_{\text{obs}})$ where $\delta_{\text{obs}}$ is the observed maximum influence.
+Computation.
+Thanks to Proposition 1, our procedure is practical and computationally convenient, allowing for application to large and varied datasets.
+While accurately estimating Extreme Value Distributions (EVDs) is a recognized and separate challenge in its own right, the maximum likelihood steps within our framework are simple and well-behaved, optimizing over only two parameters in the Gumbel case.
+The primary computational constraint instead stems from finding most influential sets — we need to approximate $\Delta^{\max}$ for the $M$ block maxima estimates.
+For computational tractability, we use an adaptive greedy algorithm (Hu et al., 2024; Kuschnig et al., 2021) with complexity $\mathcal{O}(Mk)$ and an efficient implementation based on our closed-form influence formula for sets.
 4 Experiments
-We validate our theoretical predictions and demonstrate practical utility through controlled simulations and real-world applications spanning economics, biology, and machine learning.
+In this section, we validate our theoretical predictions, investigate convergence in small samples, and demonstrate their practical relevance and utility through real-world applications spanning economics, biology, and machine learning.
 4.1 Simulation Study
+We begin with a controlled setting, where we (i) illustrate, (ii) investigate convergence for small samples, and (iii) evaluate empirical estimation.
+Illustration.
 Figure 2 illustrates our approach on a simple linear regression with one moderately influential point due to high leverage.
@@
 Panel A depicts observations, estimated regression lines with and without the influential point, and conditional significance regions at the 10, 5, and 1% levels (dotted lines).
-Panel B shows the extreme value analysis: a histogram of block maxima, fitted Gumbel distribution with (solid) and without (dashed) bias correction, and the resulting $p$-value for the observation of interest.
+Panel B illustrates the extreme value analysis: a histogram of block maxima in the background, fitted Gumbel distributions with (solid) and without (dashed) bias correction, and the resulting $p$-value for the observation of interest.
 4.1.1 Convergence to Extreme Value Distributions
-We verify that maximal influence converges to the predicted extreme value distributions.
-Across for scenarios combining standard Normal and $t(5)$ distributions for $X,R,$ we simulate $1,000$ datasets of sizes $N\in\{100,200,500,1000,2000\}.$
-Table 1 shows inverse shape estimates $\hat{\xi}^{-1}$ for the setup with $N=100.$ As predicted by theory, the Normal-Normal case yields clear Gumbel behavior ($\xi^{-1}=0$) for all sample sizes (see Table A1).
-The heavy-tailed cases exhibit predicted Fréchet behavior, with shape estimates matching theoretical predictions ($\xi^{-1}=0.2$), confirming the the finite sample applicability of Theorem 1.
-Particularly the convergence of the t(5)t(5)–Normal case has slower convergence rates to the limiting distribution due to the instability of $D^{-1}_{\mathbb{S}}$ for small samples (see Figure A2).
-Table 1: Inverse Shape Estimates for Different Distributions
-Mean Std.Dev. Q25 Median Q75
-Normal–Normal $0.0416$ $0.0319$ 0.0104-0.0104 $0.0425$ $0.0954$
-t(5)t(5)–Normal $0.1432$ $0.0363$ $0.0814$ $0.1431$ $0.2028$
-Normal–$t(5)$ $0.1706$ $0.0362$ $0.1110$ $0.1697$ $0.2316$
-$t(5)t(5)$ $0.2120$ $0.0384$ $0.1497$ $0.2112$ $0.2758$
-$1,000$ repetitions for $N=100$ observations.
-4.1.2 Location and Scale Estimation
-We evaluate whether estimation of location and scale parameters with block maximum MLE can accurately capture the true distribution.
-(Results are provided in Figure A3.) The finite sample adjusted location parameter works well and shows convergence to the true value, while the scale exhibit slight downward bias that is consistent with known limitations of MLE for EVDs (Dombry & Ferreira, 2019).
-Panel (a) in Figure A3 summarises the effectiveness of the MLE in the Gumbel case, allowing us to conduct hypothesis tests.
+Next, we verify that maximal influence converges to the predicted extreme value distributions.
+We consider four scenarios based on combinations of the standard Normal and $t(5)$ distributions for $X,R.$ For each scenario, we simulate $1,000$ datasets of sizes between $N=20$ and $N=1000,$ and compare the empirical parameter estimates with the theoretical prediction.
+Overall, we find rapid convergence, implying that our theoretical predictions are applicable with small samples.
+Figure 3 shows convergence of the scenarios to the predictions, which are indicated by dashed horizontal lines.
+(Details are provided in Table A1 of the Appendix.) All scenarios reliably converge for moderate sample sizes.
+The Normal-Normal scenario is insignificantly different from Gumbel behavior ($\xi^{-1}=0$) for $N\geqslant 50,$ and the heavy-tailed cases also exhibit the predicted Fréchet behavior ($\xi^{-1}=0.2$).
+Notably, the t(5)t(5)–Normal case converges at slower rates, likely due to the relative instability of the inverse $\left(\mathbf{X}^{\prime}\mathbf{X}\right)^{-1}_{\mathbb{S}}$ in small samples.
+Overall, the simulation results support the applicability of Theorem 1 in small samples.
+Figure 3: Convergence of empirical samples to the limiting distribution with increasing sample sizes ($N\in\{20,30,50,75,100,150,200,300,500,1000\},$ note the non-linear scale) for four different scenarios.
+Dots indicate estimated means from $1000$ repetitions, while the shaded area indicates $\pm 1\mathrm{SD}$.
+The Normal–Normal and the $t(5)t(5)$ cases quickly converge to the theoretical predictions of $\xi^{-1}=0$ and $\xi^{-1}=0.2,$ while the mixed scenarios converge marginally slower.
+Location and Scale Estimation.
+Next, we evaluate whether block maximum MLE accurately captures the location and scale parameters for empirical testing.
+Results are provided in Figure A1 of the Appendix.
+We find that bias-corrected estimation of the location parameter works well, while the scale estimate is consistent but exhibits a minor downward bias that disappears asymptotically (consistent with known limitations of the MLE; see Dombry and Ferreira, 2019).
+For our goal of hypothesis testing, the overall distribution and its quantiles are recovered effectively.
 4.2 Applications
-We investigate several real-world datasets — two applications from economics and biology, as well as four well-known machine learning benchmarks.
-4.2.1 Economic Development and Geography
-We re-examine the controversial finding that rugged terrain benefits African economies when compared to the rest of the world (Nunn, 2020).
+We investigate several real-world datasets — two applications from economics and biology, and four machine learning benchmarks — and provide the first conclusive investigation of influence.
+Economic Development and Geography.
+We re-examine the controversial finding that rugged terrain benefits African economies when compared to the rest of the world (Nunn and Puga, 2012).
 Kuschnig et al.
-(2021) identify the Seychelles, coupled with any of Rwanda, Lesotho, Eswatini, and the Comoros, as influential, removing significance of the estimate of interest.
-Our results decisively resolve this controversy.
-Table 2 reveals the Seychelles as excessively influential on $\hat{\theta}_{\text{rugged}}$, both individually ($p<0.001$) and in combination with other outliers except for Lesotho.
-This confirms the suspected confounding from the size of nations, lending statistical rigor to prior concerns.
+(2021) call the influence of the Seychelles, which remove significance of the estimate of interest when coupled with any of Rwanda, Lesotho, Eswatini, and the Comoros.
+However, lacking the statistical framework that we provide, they were not able to test whether this level of influence should be deemed excessive or not.
+We can now decisively resolve this controversy.
+Table 1 reveals the Seychelles as excessively influential on $\hat{\theta}_{\text{rugged}}$, both individually ($p<0.001$) and in combination with other outliers, except for Lesotho.
+This substantiates the suspected confounding from the size of nations (Kuschnig et al., 2021), lending statistical rigor to prior concerns and calling into question the differential relation between ruggedness and income in Africa.
 Table 1: Influence of Ruggedness on log(GDP per capita in 2000)
@@
 Seychelles + Rwanda $0.070$ $0.028$ $0.006$ $0.001$
-Seychelles + Eswatini $0.077$ $0.020$ $0.004$ $1e^{-16}$
+Seychelles + Eswatini $0.091$ $0.027$ $0.006$ $1e^{-16}$
 Seychelles + Comoros $0.061$ $0.028$ $0.006$ $0.004$
-4.2.2 Sparrow Morphology — Head and Beak Size
-We analyze the relation between head and tarsus length in saltmarsh sparrows, based on measurements of $N=1295$ sparrows with known outliers (Gjerdrum et al., 2008; Zuur et al., 2010).
+Sparrow Morphology — Big Heads and Beaks.
+We analyze the relation between head and tarsus length in saltmarsh sparrows, based on measurements of $N=1,295$ sparrows with known outliers (Gjerdrum et al., 2008; Zuur et al., 2010).
 The baseline regression yields $\hat{\theta}=0.011$ with a standard error of $(.030),$ implying a relation that is statistically indistinguishable from zero.
@@
 An additional data point further moves the estimate to $0.288(.032).$ These extreme impacts from a vanishing fraction of the sample are deemed excessive by our approach at any conventional significance level (both p<0.001p<0.001).
-[^3] A possible explanation for this excessive influence are data entry errors: The first observation (an outlier in both head and tarsus size) may have the two (adjacent) features mixed up — when swapped, they would fit well into overall averages.
+[^5] One possible explanation for this excessive influence are data entry errors: The first observation (an outlier in both head and tarsus size) may have the two (adjacent) features mixed up — when swapped, they would fit well into overall averages.
 The second observation (an outlier in one feature) stands out with both values being equal up to the one significant digit.
-4.2.3 Machine Learning Benchmarks
+Machine Learning Benchmarks.
 We apply our framework to four widely-used regression benchmarks: Law School, Adult Income, Boston Housing, and Communities & Crime.
@@
 We consider two sets: $77$ data points that move the estimate from 0.0412(.0144)-0.0412\ (.0144) to 0.1117(.0159),0.1117\ (.0159), creating a significant estimate with flipped sign, and $17$ data points that reduce the estimate to 0.0223(.0097).-0.0223\ (.0097).
-Our approach indicates that the influence larger set’s influence falls within expected variation, while the smaller set exhibits statistically excessive influence ($p=0.019$).
+Our approach indicates that the larger set’s influence falls within expected variation, while the smaller set exhibits statistically excessive influence ($p=0.019$).
   • Adult Income ($N=32,561$): We investigate the top 1% most influential sets ($325$ points) that shift the ‘Male’ indicator from $\hat{\theta}=0.062,$ either raising it to $0.0992$ or decreasing it to $0.0214.$ Despite these considerable shifts from a small fraction of the data, neither is deemed excessively influential by our approach.
@@
 The baseline (highly significant) coefficient 0.1080(.0329)-0.1080\ (.0329) is rendered insignificant at 0.0352(.0556)-0.0352\ (.0556) after excluding just $6$ observations.
-In this case, the underlying EVD is Frechét with inverse shape $\xi^{-1}=0.29$ due to the heavy tail of the crime variable.
+In this case, the underlying EVD is Fréchet with inverse shape $\xi^{-1}=0.29$ due to the heavy tail of the crime variable.
 The set’s influence is highly significant ($p=0.001$), indicating excessive influence.
@@
 The complete set is not extreme, as the points cancel each other out.
-After exclusion, the first subset of two increases the coefficient by more than 22%, which is deemed excessive $p<0.001.$ When re-estimating after their exclusion, the second set decreases the estimate by more than 10% and is deemed excessive at the 5% level ($p=0.014.$ (See Table A2 for details.)
+After exclusion, the first subset of two increases the coefficient by more than 22%, which is deemed excessive $p<0.001.$ When re-estimating after their exclusion, the second set decreases the estimate by more than 10% and is deemed excessive at the 5% level ($p=0.014$).
+(See Table A2 for details.)
 5 Discussion
-Our analysis provides the first rigorous statistical framework for assessing when most influential sets represent genuine problems rather than natural sampling variation.
-By establishing that maximal influence follows predictable extreme value distributions, we enable practitioners to move beyond ad-hoc rules and domain-specific judgment.
-The key insight is that maximal influence fundamentally depends on the nature of the sets considered and the tail behavior of the underlying data.
-This work addresses a critical gap in interpretable machine learning, where theoretical foundations for influential set analysis have been lacking.
-We show that maximal influence depends on the tail properties of the underlying data—when tails are heavy, influence patterns become unpredictable, requiring robust estimation methods or acceptance of the inherent instability.
-For distributions with moderate tail behavior, our results provide tight theoretical bounds.
-Our approach integrates naturally with the literature on influential set selection (Broderick et al., 2021; Fisher et al., 2023; Hu et al., 2024), which has lacked conclusive theoretical guidance.
-While influential sets share connections with established diagnostics such as Cook’s distance and leverage scores, or methods such as robust regression (Huber & Ronchetti, 2009), they fill a unique role.
-By directly relating to quantities of interest, they allow practitioners to discover and analyze sets of data that deviate from dominant patterns and yield insights that other methods obscure (see, e.g., Kuschnig et al., 2021, for comparisons).
-Our results provide the theoretical foundations for analyzing these influential sets.
-[^4] We can also clarify the applicability of the common $2\sqrt{N}$ threshold for coefficient influence (Belsley et al., 1980).
-While imprecise for most influential observations — missing the selection procedure that necessitates extreme value theory — it proves asymptotically accurate for randomly selected observations.
-5.1 Limitations and Future Work
-Our analysis operates within linear regression — a foundational setting for theory and modern ML, but limited to contexts where interpretability is valued (Rudin, 2019; Roscher et al., 2020).
-While we focus on regression coefficients, an extension predictions follows trivially through $\hat{Y}=X\hat{\beta}$.
-Extending our results to generalized linear, tree-based, or non-parametric models, requires further developments.
-The asymptotic arguments presented here provide important insights, but come with inherent limitations.
-Our analysis leverages independence between features and the residuals, which can be restrictive in practice where dependence affects influence patterns.
-While we address these concerns in controlled settings, the gap between theory and finite-sample behavior warrants investigation.
-Several methodological improvements could strengthen practical performance.
-Estimation of extreme value parameters could be enhanced through setting-specific methods and improved bias correction (Dombry & Ferreira, 2019).
-The efficient selection of most influential sets themselves remains an open challenge (Hu et al., 2024; Huang et al., 2025) with direct implications for our procedure.
-5.2 Broader Implications
-This work enables more reliable and transparent decision-making across domains where linear models remain the method of choice.
+We develop the first rigorous statistical framework for assessing when most influential sets represent genuine problems rather than natural sampling variation.
+By deriving the extreme value distribution of maximal influence, we allow practitioners to replace ad-hoc sensitivity checks with principled statistical decision rules.
+Our key insight is how maximal influence fundamentally depends on set size and tail behavior.
+For constant-size sets with polynomial tails, maximal influence follows a heavy-tailed Fréchet distribution, implying that extreme influence can be arbitrarily large.
+For growing sets or exponential-tailed data, maximal influence converges to a well-behaved Gumbel distribution.
+These results address a critical gap in interpretable machine learning.
+While recent work has developed methods to identify influential sets (Broderick et al., 2023; Freund and Hopkins, 2023; Hu et al., 2024), no formal theory existed to determine when their influence is excessive.
+Our framework provides the long-missing theoretical foundation that enables rigorous statistical inference for influential observations and sets (first discussed by Cook, 1979).
+We can clarify the applicability of heuristics that are commonly used for identifying excessive influence, and provide conclusive answers when excessively influential sets are suspected (such as the Seychelles in the ruggedness example; see Kuschnig et al., 2021).
+The $2/\sqrt{N}$ threshold for coefficient influence (Belsley et al., 1980), e.g., is asymptotically accurate for randomly selected observations, but is too restrictive for most influential observations, where the selection procedure necessitates extreme value theory.
+Practical Recommendations.
+In general, most influential sets hold valuable information for inference.
+Our test is deliberately conservative, controlling Type I errors (false claims of excessive influence) at the cost of some Type II errors (failing to detect truly excessive influence).
+This reflects our view that influential sets are a natural feature of data and not a problem to be eliminated.
+When our test identifies an excessively influential set, however, we recommend:
+  1. 1.
+    Investigate mechanism.
+Document the set and investigate why it differs — it may convey genuine heterogeneity, data quality issues, unobserved confounding, or important edge cases not addressed by the model.
+  2. 2.
+    Handle appropriately.
+We argue that an excessively influential set warrants separate analysis; exclusion can be considered if it reflects measurement error or outliers that are irrelevant for the pattern of interest.
+  3. 3.
+    Report transparently.
+State the set, decision, and test outcome; if conclusions hinge on the set, report both and discuss why.
+We advise against trimming or winsorizing to force alignment with the remaining data; these transformations create artificial data that may obscure and distort rather than illuminate underlying relations.
+Limitations and Future Work.
+Our analysis focuses on linear regression — foundational for theory and modern ML methods, but limited to contexts where interpretability is valued (Rudin, 2019).
+Extensions to generalized linear models, tree-based methods, or non-parametric estimators require further developments.
+Our asymptotic arguments leverage independence between features and residuals, which may be restrictive when dependence affects influence patterns.
+Explicitly addressing dependence between influence components through the selected set and across sets is possible through generalizations.
+While simulations show that small-sample behavior quickly converges to asymptotic predictions (even at $N=100$), further investigation of the theory-practice gap is warranted.
+Practicality of our approach hinges on estimating the extreme value null, and, hence, approximating maximum influence.
+Finite-sample estimation of tail behavior and EVD parameters can be delicate, and improving these estimators would directly sharpen $p$-values.
+Two options include leveraging domain-specific information and improved bias correction methods (Dombry and Ferreira, 2019; Oorschot and Zhou, 2020).
+Advances in finding influential sets remain an active research area (Hu et al., 2024; Huang et al., 2025), and could substantially improve block-maxima, reduce runtime, and broaden applicability.
+Broader Implications.
+Our framework enables more reliable decision-making across domains where linear models remain the method of choice.
 Principled tools for understanding data points that drive model behavior are crucial for building trustworthy systems.
-Applications span fairness assessments, where influential subsets can reveal algorithmic bias, to causal inference settings, including randomized controlled trials quasi-experimental econometric analyses where small data subset can fundamentally alter estimates.
-Importantly, we reframe influence as a natural and informative feature of data that requires appropriate treatment rather than a probem to be fixed.
+Applications span fairness assessments — where influential sets can reveal algorithmic bias — to causal inference settings, such as randomized controlled trials or quasi-experimental econometric analyses where small data subsets can fundamentally alter estimates.
+Importantly, we reframe influence as a natural feature of data requiring appropriate treatment rather than a problem to be fixed.
 Influential sets can represent genuine heterogeneity or important edge cases that should inform model development.
-This perspective emables more nuanced approaches to data processing and model validation, where more information is preserved and assessed through principled statistical inference rather than discarded based on rules of thumb.
+This perspective enables more nuanced approaches to data analysis, where information is preserved and assessed through principled statistical inference rather than discarded based on rules of thumb.
 6 Conclusion
@@
 Reproducibility Statement
-Theoretical results are elaborated upon in the Appendix, where proofs are elaborated upon via several Lemmata.
-Datasets can be obtained from the cited sources, and code that generates the results will be made available.
+Proofs are detailed in the Appendix, datasets are from the cited sources, and code to reproduce results are available at https://github.com/konradld/testingMIS.
 Statement on LLM Use
-Large language models were used to (i) aid and polish writing, (ii) retrieve and discover related work, and (iii) check results for apparent mistakes.
-Appendix A1 Exact Influence Formulas
-Proposition 2.
-Consider the influence in Equation 3. We can simplify this expression to
-$\Delta\left(\{i\}\right)=\frac{x_{i}r_{i}}{\sum_{n\neq i}x_{n}^{2}}.$
-Proof.
-Let $A_{i}\mathrel{\raisebox{0.0pt}{:}}=x_{i}^{2}$ and $D\coloneqq\sum_{n}^{N}X_{n}^{2},$ such that $D_{-\{i\}}=\sum_{n\neq i}^{N}x_{n}^{2}.$ Then
-$\Delta\left(\{i\}\right)=\frac{\frac{\sqrt{A_{i}}r_{i}}{(A_{i}+D_{-\{i\}})}}{1-\frac{A_{i}}{(A_{i}+D_{-\{i\}})}}$ $=\frac{(A_{i}^{1.5}\cdot r_{i}+A_{i}^{0.5}\cdot r_{i}\cdot D_{-\{i\}})}{D_{-\{i\}}\left(A_{i}+D_{-\{i\}}\right)}$
-$=\frac{r_{i}\cdot x_{i}\cdot\left(A_{i}+D_{-\{i\}}\right)}{D_{-\{i\}}\cdot\left(A_{i}+D_{-\{i\}}\right)}=\frac{r_{i}\cdot x_{i}}{\sum_{n\neq i}^{N}x_{n}^{2}}.$
-A1.1 Recursion
-It helps to know the influence on the residual, leverage, and full hat matrix:
-$(r_{i})_{-\{j\}}$ $=r_{i}+x_{i}\,\Delta\left(\{j\}\right),$
-$(h_{i})_{-\{j\}}$ $=x_{i}^{2}/\sum_{n\neq j}x_{n}^{2},$
-$(h_{ij})_{-\{k\}}$ $=h_{ij}+\frac{h_{ik}h_{kj}}{1-h_{k}}.$
-Proposition 3.
-Let ${\mathbb{S}}=\left\{1,\dots,K\right\}.$ Then we can recursively define $\Delta\left({\mathbb{S}}\right)$ as
-$\Delta\left({\mathbb{S}}\right)$ $=\hat{\beta}-\hat{\beta}_{-{\mathbb{S}}}$
-$=\Delta\left(\{1\}\right)+\Delta\left(\{2\}\right)_{-\{1\}}+\dots+\Delta\left(\{K\}\right)_{-{\mathbb{S}}\setminus K}.$
-Proof.
-Notice that
-[…]
-[…]
-$=\Delta\left(\left\{j\right\}\right)+\Delta\left(\left\{i\right\}\right)_{-\{j\}}.$
-Equivalent results trivially hold for larger sets. ∎
-Now, we can provide the full details for the induction step in the proof of Proposition 1:
-Proof.
-$\Delta\left({\mathbb{S}}\right)=\frac{\sum_{k=1}^{K-1}x_{k}r_{k}}{D_{-{\mathbb{S}}}+x_{k}^{2}}+\frac{x_{k}r_{k}}{D_{-{\mathbb{S}}}}+\frac{x_{k}^{2}\sum_{k=1}^{K-1}x_{k}r_{k}}{D_{-{\mathbb{S}}}\left(D_{-{\mathbb{S}}}+x_{k}^{2}\right)}=\frac{\sum_{k=1}^{K}x_{k}r_{k}}{D_{-{\mathbb{S}}}},$
-and
-$\Delta\left({\mathbb{S}}\right)$ $=\frac{\sum_{k=1}^{K}x_{k}r_{k}}{D_{-{\mathbb{S}}}+x_{K+1}^{2}}+\frac{x_{K+1}r_{K+1}}{D_{-{\mathbb{S}}}}+\frac{x_{K+1}^{2}\sum_{k=1}^{K}x_{k}r_{k}}{D_{-{\mathbb{S}}}\left(D_{-{\mathbb{S}}}+x_{K+1}^{2}\right)}$
-$=\frac{\sum_{i=k}^{K}x_{k}r_{k}\left(D_{-{\mathbb{S}}}+x_{K+1}^{2}\right)}{D_{-{\mathbb{S}}}\left(D_{-{\mathbb{S}}}+x_{K+1}^{2}\right)}+\frac{x_{K+1}r_{K+1}}{D_{-{\mathbb{S}}}}=\frac{\sum_{k=1}^{K+1}x_{k}r_{k}}{D_{-{\mathbb{S}}}}.$
+Large language models were used to (i) aid and polish writing, (ii) discover and retrieve related work, (iii) check results for apparent mistakes.
 Appendix A1 Lemma for the Inverse Sum of Squares
@@
   3. 3.
-    $\operatorname{Var}(X_{1}^{2})=\sigma^{2}>0$ (non-degenerate variance of squares)
+    $\mathrm{Var}(X_{1}^{2})=\sigma^{2}>0$ (non-degenerate variance of squares)
 Define $S_{n}=\sum_{i=1}^{n}X_{i}^{2}$ and $Y_{n}=S_{n}^{-1}$. Then $Y_{n}$ is asymptotically normal with:
@@
 $\sqrt{n}\left(\bar{X}_{n}^{(2)}-\mu\right)\xrightarrow{d}\mathcal{N}(0,\sigma^{2}),$
-where $\mu=\mathbb{E}[X_{1}^{2}]$ and $\sigma^{2}=\operatorname{Var}(X_{1}^{2})$ (finite by $\mathbb{E}[X_{1}^{4}]<\infty$).
+where $\mu=\mathbb{E}[X_{1}^{2}]$ and $\sigma^{2}=\mathrm{Var}(X_{1}^{2})$ (finite by $\mathbb{E}[X_{1}^{4}]<\infty$).
 Consider the transformation $g(x)=x^{-1}$, which is differentiable at $x=\mu>0$ with derivative $g^{\prime}(x)=-x^{-2}$. The Delta Method gives:
@@
 For notational simplicity let $S\coloneqq\sum_{i\in{\mathbb{S}}}X_{i}\cdot R_{i}$ and $T\coloneqq D^{-1}_{-{\mathbb{S}}},$ where $D=\sum_{n}X_{n}^{2}$.
-It holds for any realization $S=s\in\mathbb{R}$ and $T=t\in\mathbb{R^{+}}$.
+This holds for any realization $S=s\in\mathbb{R}$ and $T=t\in\mathbb{R}^{+}$.
 Further, let $\mathrm{MDA}(H)$ denote the maximum domain of attraction of an EVD $H$ where we write $Z\in\mathrm{MDA}(H)$.
 We specifically denote the Fréchet as $\Phi_{\alpha}$ and the Gumbel as $\Lambda.$ We are interested in the EVD of $\Delta=S\cdot T.$
-A3.1 If $S\in\mathrm{MDA}(\Lambda)$
 Lemma 2.
-Let $T\in\mathrm{MDA}(\Lambda)$ and $S\in\mathrm{MDA}(\Phi_{a})$ with tail-coefficient $a>0$ and $S$ and $T$ being independent, then $\Delta\left({\mathbb{S}}\right)=S\cdot T\in\mathrm{MDA}(\Phi_{a})$ .
+Let $S\in\mathrm{MDA}(\Lambda)$ and $T\in\mathrm{MDA}(\Phi_{a})$ with tail-coefficient $a>0$ and $S$ and $T$ being independent, then $\Delta\left({\mathbb{S}}\right)=S\cdot T\in\mathrm{MDA}(\Phi_{a}).$
 Proof.
-Recall that for Gumbel tails ($S$) the survival function decays double-exponentially, i.e.,
-$\mathbb{P}(S>s)\sim\exp\left(-\exp\left(\frac{s-\mu}{\beta}\right)\right)\quad\text{as }s\to\infty,$
+Recall that for Gumbel tails ($S$) the survival function decays faster than any polynomial, i.e.,
+$\mathbb{P}(S>s)\sim\exp\left(-\frac{s-\mu}{\beta}\right)\quad\text{as }s\to\infty,$
 while for the Fréchet tails ($T$) the survival function is regularly varying with index a-a, i.e.,
@@
 $f_{T}(t)\sim at^{-a-1}L_{T}(t)\quad\text{as }t\to\infty.$
-We are interested in the EVD of $\Delta$, i.e., $\mathbb{P}(\Delta>\delta)$. Since $\Delta=S\cdot T$ and $S$, $T$ are independent by construction:
+We are interested in the EVD of $\Delta$, i.e., $\mathbb{P}(\Delta>\delta)$. Since $\Delta=S\cdot T$ and $S$, $T$ are independent by assumption:
 $\mathbb{P}(\Delta>\delta)=\mathbb{P}(S\,T>\delta)=\int_{\mathbb{R}^{+}}\mathbb{P}(S>\delta/t)f_{T}(t)\,\mathrm{d}t.$
@@
 As $\delta\to\infty$, by Lemma 4 in Appendix A3, we obtain
-$\int_{0}^{\delta/M}\mathbb{P}(S>u)u^{a-1}L_{T}\left(\frac{\delta}{u}\right)\,\mathrm{du}\sim L_{T}(\delta)\int_{0}^{\infty}\mathbb{P}(S>u)u^{a-1}\,\mathrm{du}.$ (A1)
+$\int_{0}^{\delta/M}\mathbb{P}(S>u)u^{a-1}L_{T}\left(\frac{\delta}{u}\right)\,\mathrm{d}u\sim L_{T}(\delta)\int_{0}^{\infty}\mathbb{P}(S>u)u^{a-1}\,\mathrm{d}u.$ (A2.1)
 The integral converges because:
@@
 $\mathbb{P}(\Delta>\delta)\sim a\delta^{-a}L_{T}(\delta)C(a,S)=\delta^{-a}\left(aC(a,S)L_{T}(\delta)\right).$
-The term in parentheses is slowly varying in $\delta$ since $L_{T}(\delta)$ is slowly varying which concludes the proof. ∎
-To summarize, the survival function $\mathbb{P}(\Delta>\delta)$ is regularly varying with index a,-a, and therefore, $\Delta$ has Fréchet tails with tail-coefficient $a.$
+The term in parentheses is slowly varying in $\delta$ since $L_{T}(\delta)$ is slowly varying.
+Thus, the survival function $\mathbb{P}(\Delta>\delta)$ is regularly varying with index a,-a, and therefore, $\Delta$ has Fréchet tails with tail-coefficient $a$, which concludes the proof. ∎
 Corollary 2.
@@
 Further, if $S\in\mathrm{MDA}(E)$ for some EVD $E$, it holds that
-$\Delta\left({\mathbb{S}}\right)\mid X_{-{\mathbb{S}}}\in\mathrm{MDA}(E),$
+$\Delta\left({\mathbb{S}}\right)\mid X_{-{\mathbb{S}}}\in\mathrm{MDA}(E).$
 Appendix A3 Lemma for Asymptotic Equivalence
@@
     $\left|\frac{L_{T}(\delta/u)}{L_{T}(\delta)}\right|\leqslant C_{\delta}u^{\delta}\quad\text{for all large }\delta,u\geqslant 1.$
-    Choose $\delta$ small enough such that $k=a-1+\delta>0,$ and
+    Choose $\delta$ such that $k=a-1+\delta>0,$ and
     $\int_{1}^{\infty}\mathbb{P}(S>u)u^{k}\,\mathrm{d}u<\infty,$
@@
 $R(\delta)=\int_{\delta/M}^{\infty}\mathbb{P}(S>u)u^{a-1}L_{T}\left(\frac{\delta}{u}\right)\,\mathrm{d}u.$
-  • For $u\geqslant\delta/M$, we have $\delta/u\leqslant M$ s.t. is bounded on compact sets: $L_{T}(\delta/u)\leqslant C_{M}.$
+  • For $u\geqslant\delta/M$, we have $\delta/u\leqslant M,$ which is bounded on compact sets: $L_{T}(\delta/u)\leqslant C_{M}.$
   • By the Gumbel tail properties, there exist a $\theta>0$ s.t. $\mathbb{P}(S>u)\leqslant e^{-u^{\theta}}$ for large $u.$ Thus
@@
 $=$ $\int_{0}^{\infty}\mathbb{P}(S>u)u^{a-1}\,\mathrm{d}u.$
-Appendix A5 Estimation
-Figure A1: Visualization of Table 1.
-Most notable is the regime change between thin and polynomial tails.
-While for the Frechét cases the average MLE is statistically insignificantly different form another, they are all individually different from the Gumbel case.
-While the finite sample $D^{-1}_{-{\mathbb{S}}}$ has thicker than Normal tails, acting as a regularizer on $\Delta^{\max}$ and thereby resulting in slight underestimation of the theoretically suggested $\xi^{-1}=0.20$.
-Table A1: Inverse Shape Estimates for Different Distributions and Sample Sizes
+Appendix A4 Auxiliary Material
 NN Distribution Mean Std.Dev. Q25 Median Q75
-100 Normal–Normal $0.0416$ $0.0319$ 0.0104-0.0104 $0.0425$ $0.0954$
-t(5)t(5)–Normal $0.1432$ $0.0363$ $0.0814$ $0.1431$ $0.2028$
-Normal–$t(5)$ $0.1706$ $0.0362$ $0.1110$ $0.1697$ $0.2316$
-$t(5)t(5)$ $0.2120$ $0.0384$ $0.1497$ $0.2112$ $0.2758$
-200 Normal–Normal $0.0267$ $0.0318$ 0.0249-0.0249 $0.0272$ $0.0786$
-t(5)t(5)–Normal $0.1394$ $0.0361$ $0.0774$ $0.1400$ $0.1967$
-Normal–$t(5)$ $0.1666$ $0.0353$ $0.1091$ $0.1670$ $0.2241$
-$t(5)t(5)$ $0.2033$ $0.0374$ $0.1406$ $0.2049$ $0.2627$
-500 Normal–Normal $0.0147$ $0.0316$ 0.0358-0.0358 $0.0159$ $0.0635$
-t(5)t(5)–Normal $0.1509$ $0.0387$ $0.0854$ $0.1528$ $0.2133$
-Normal–$t(5)$ $0.1692$ $0.0367$ $0.1076$ $0.1696$ $0.2316$
-$t(5)t(5)$ $0.2088$ $0.0382$ $0.1437$ $0.2096$ $0.2685$
-1000 Normal–Normal $0.0101$ $0.0333$ 0.0451-0.0451 $0.0100$ $0.0641$
-t(5)t(5)–Normal $0.1596$ $0.0378$ $0.0941$ $0.1603$ $0.2215$
-Normal–$t(5)$ $0.1726$ $0.0360$ $0.1137$ $0.1727$ $0.2333$
-$t(5)t(5)$ $0.2101$ $0.0374$ $0.1495$ $0.2123$ $0.2689$
-2000 Normal–Normal $0.0106$ $0.0327$ 0.0431-0.0431 $0.0114$ $0.0626$
-t(5)t(5)–Normal $0.1731$ $0.0371$ $0.1113$ $0.1740$ $0.2323$
-Normal–$t(5)$ $0.1759$ $0.0359$ $0.1185$ $0.1762$ $0.2387$
-$t(5)t(5)$ $0.2150$ $0.0375$ $0.1521$ $0.2139$ $0.2763$
-All results based on 1000 repetitions.
-Figure A2: Convergence of cases to the limiting distribution, visualizing Table A1.
-Solid lines are estimated means from $1000$ repetitions while the shaded area is $\pm 1SD$.
-The Normal–Normal case as well as the $t(5)t(5)$ case converge quickly to the theoretical $\xi^{-1}=0$ and $\xi^{-1}=0.2$ respectively.
-The convergence rate of the mixed cases converge slower yet reasonably fast.
-In particular, the t(5)t(5)–Normal case is visibly requiring more observations to stabilize $D^{-1}_{\mathbb{S}}$.
-In general this finding supports the relevance and applicability of our limiting results to finite samples.
+20 Normal–Normal $0.1385$ $0.2541$ 0.0206-0.0206 $0.1293$ $0.2902$
+t(5)t(5)–Normal $0.2417$ $0.2865$ $0.0863$ $0.2329$ $0.3994$
+Normal–$t(5)$ $0.2529$ $0.2479$ $0.0910$ $0.2505$ $0.4089$
+$t(5)t(5)$ $0.3271$ $0.2719$ $0.1601$ $0.3144$ $0.4922$
+30 Normal–Normal $0.1021$ $0.1803$ 0.0079-0.0079 $0.1101$ $0.2291$
+t(5)t(5)–Normal $0.1989$ $0.1845$ $0.0917$ $0.1994$ $0.3067$
+Normal–$t(5)$ $0.2023$ $0.1846$ $0.0817$ $0.2101$ $0.3225$
+$t(5)t(5)$ $0.2670$ $0.1912$ $0.1453$ $0.2668$ $0.3961$
+50 Normal–Normal $0.0668$ $0.1255$ 0.0153-0.0153 $0.0631$ $0.1466$
+t(5)t(5)–Normal $0.1708$ $0.1304$ $0.0890$ $0.1701$ $0.2532$
+Normal–$t(5)$ $0.1843$ $0.1286$ $0.1015$ $0.1797$ $0.2704$
+$t(5)t(5)$ $0.2307$ $0.1363$ $0.1420$ $0.2305$ $0.3203$
+75 Normal–Normal $0.0474$ $0.0960$ 0.0163-0.0163 $0.0485$ $0.1176$
+t(5)t(5)–Normal $0.1500$ $0.1006$ $0.0821$ $0.1469$ $0.2173$
+Normal–$t(5)$ $0.1735$ $0.1057$ $0.1037$ $0.1749$ $0.2434$
+$t(5)t(5)$ $0.2214$ $0.1046$ $0.1561$ $0.2213$ $0.2885$
+100 Normal–Normal $0.0422$ $0.0788$ 0.0115-0.0115 $0.0429$ $0.0959$
+t(5)t(5)–Normal $0.1389$ $0.0846$ $0.0790$ $0.1369$ $0.1948$
+Normal–$t(5)$ $0.1721$ $0.0865$ $0.1183$ $0.1731$ $0.2291$
+$t(5)t(5)$ $0.2083$ $0.0876$ $0.1492$ $0.2056$ $0.2639$
+200 Normal–Normal $0.0253$ $0.0506$ 0.0075-0.0075 $0.0253$ $0.0582$
+t(5)t(5)–Normal $0.1418$ $0.0587$ $0.1020$ $0.1403$ $0.1833$
+Normal–$t(5)$ $0.1655$ $0.0576$ $0.1272$ $0.1632$ $0.2013$
+$t(5)t(5)$ $0.2072$ $0.0595$ $0.1691$ $0.2054$ $0.2450$
+300 Normal–Normal $0.0169$ $0.0406$ 0.0096-0.0096 $0.0180$ $0.0455$
+t(5)t(5)–Normal $0.1423$ $0.0461$ $0.1122$ $0.1446$ $0.1731$
+Normal–$t(5)$ $0.1705$ $0.0463$ $0.1377$ $0.1695$ $0.2006$
+$t(5)t(5)$ $0.2082$ $0.0511$ $0.1757$ $0.2082$ $0.2406$
+500 Normal–Normal $0.0157$ $0.0319$ 0.0049-0.0049 $0.0154$ $0.0381$
+t(5)t(5)–Normal $0.1488$ $0.0378$ $0.1245$ $0.1499$ $0.1739$
+Normal–$t(5)$ $0.1667$ $0.0354$ $0.1419$ $0.1662$ $0.1903$
+$t(5)t(5)$ $0.2087$ $0.0385$ $0.1833$ $0.2085$ $0.2323$
+1000 Normal–Normal $0.0103$ $0.0220$ 0.0042-0.0042 $0.0109$ $0.0243$
+t(5)t(5)–Normal $0.1591$ $0.0287$ $0.1408$ $0.1589$ $0.1771$
+Normal–$t(5)$ $0.1734$ $0.0252$ $0.1565$ $0.1743$ $0.1905$
+$t(5)t(5)$ $0.2105$ $0.0262$ $0.1938$ $0.2101$ $0.2273$
+Table A1: Inverse shape estimates for different distributions and sample sizes.
+They indicate fast convergence to asymptotic predictions ($0.0$ for the Normal-Normal, and $0.2$ for other cases) — for all settings, the predicted value is contained between the 25th and 75th quantile.
+Estimates are based on 1,000 repetitions.
 Figure A1: Simulation exercise for the performance of the simple MLE based on block maxima, correcting for block size.
-While the corrected location parameter $\hat{\tilde{a}}$ is close to unbiased, the scale parameter $\hat{b}$ suffers some downward bias using simple block maxima, which is in line with Dombry & Ferreira (2019).
-However, for practical purposes the block maxima is expected to be fitting reasonably well, as visible in panel (a).
-Appendix A6 Auxiliary Results for Case Studies
-As mentioned in the main text, Table A2 summarizes the results for testing the preselected set and its subsets for significant influence of the percent of black population on the violent crimes committed per population.
-Table A2: Influence of % Black Population on Violent Crimes
+While the corrected location parameter $\hat{\tilde{a}}$ is close to unbiased, the scale parameter $\hat{b}$ suffers some downward bias using simple block maxima, which is in line with Dombry and Ferreira (2019).
+However, for practical purposes the block maxima are expected to be fitting reasonably well, as visible in panel (a).
 Set Composition Set Size $\Delta({\mathbb{S}})$ $\hat{\tilde{a}}$ $\hat{b}$ $p$-value
@@
 1st partial 22 $0.0456$ $0.0050$ $0.0021$ $7.62e^{-7}$
-2nd Partial after excl. 1st 22 0.0241-0.0241 $0.0051$ $0.0022$ $0.0141$
+2nd after excl. 1st 22 0.0241-0.0241 $0.0051$ $0.0022$ $0.0141$
+Table A2: Influence of % Black Population on Violent Crimes.
+The table summarizes the results for testing the preselected set and its subsets for significant influence of the percent of black population on the violent crimes committed per population.